### v1

# A Reproducible Framework for Neural Topic Modeling in Focus Group Analysis

Abstract
Focus group discussions generate rich qualitative data but their analysis traditionally relies on labor-intensive manual coding that limits scalability and reproducibility. We present a rigorous, reproducible computational framework for applying neural topic modeling to focus group transcripts, addressing fundamental methodological challenges: hyperparameter sensitivity, model stability, and validation of interpretability. Using BERTopic applied to ten focus groups exploring HPV vaccine perceptions in Tunisia (1,076 utterances), we conducted systematic evaluation across 27 hyperparameter configurations, assessed stability through bootstrap resampling with 30 replicates per configuration, and validated interpretability through formal human evaluation by three domain experts. Our analysis demonstrates substantial sensitivity to hyperparameter choices and reveals that metric selection for stability assessment must align with analytical goals. A hierarchical merging strategy (extracting fine-grained topics for stability then consolidating for interpretability) effectively navigates the stability-coherence tradeoff, achieving coherence of 0.558 compared to 0.539 for direct extraction. Human validation confirmed topic quality with very good inter-rater reliability (ICC = 0.79, weighted Cohen’s kappa = 0.578). Our framework provides practical guidelines that researchers can adapt to their own qualitative research contexts. All code, data processing scripts, and evaluation protocols are publicly available to support reproduction and extension of this work.
Keywords Neural Topic Modeling $\cdot$ BERTopic $\cdot$ Focus Group Analysis $\cdot$ Model Validation $\cdot$ Reproducibility $\cdot$ Hyperparameter Optimization $\cdot$ Bootstrap Resampling $\cdot$ Text Mining
1 Introduction
Topic modeling has emerged as a powerful computational approach for discovering latent thematic structures in large text corpora. Originally developed for document organization and information retrieval, topic modeling techniques have found increasing application in public health research, where understanding discourse patterns can inform intervention strategies and policy decisions. The COVID-19 pandemic accelerated this trend, with numerous studies applying topic modeling to analyze public sentiment and concerns expressed on social media platforms [qorib_covid-19_2023, figueiredo_topic_2024, yin_sentiment_2022, lyu_covid-19_2021, boon-itt_public_2020]. Similarly, topic modeling has been deployed to characterize discussions about Human Papillomaviruss (HPV) vaccination [du_use_2020, surian_characterizing_2016, rai_detecting_2024], cancer care experiences [steele2025], and mental health support communities [carron-arthur_whats_2016]. These applications demonstrate the growing recognition of topic modeling as a valuable tool for public health surveillance and understanding health-related discourse at scale.
However, the majority of these applications focus on readily available online data sources such as Twitter, Reddit, and online forums. While social media provides unprecedented access to large-scale public discourse, it represents only one facet of health communication. Focus group discussions, a cornerstone methodology in qualitative health research, offer structured settings where participants engage in moderated conversations that elicit deeper, contextualized perspectives on health beliefs, barriers, and decision-making processes. Despite their methodological importance, focus group transcripts have received limited attention in the topic modeling literature. This gap likely reflects both the proprietary nature of focus group data and the traditional emphasis on human-coded qualitative analysis. Yet focus groups generate rich textual data that could benefit from computational approaches, particularly when research objectives include systematic synthesis across multiple sessions or geographic contexts.
The application of topic modeling to qualitative research raises important methodological challenges. Focus group transcripts are typically small, heterogeneous corpora with non-uniform conversational structure, which can push topic models outside of the conditions under which they are commonly benchmarked. In such low-data regimes, model outcomes can be highly sensitive to design choices. At the same time, topic modeling algorithms are inherently stochastic, which means that seemingly minor configuration changes can lead to very different topic structures which raises questions about the reproducibility and stability of results, especially when findings inform public health interventions. Recent methodological reviews have highlighted widespread issues with reproducibility in topic modeling research [hagg2022] and emphasized the multidimensional nature of model quality [hosseiny2023].
This paper presents a rigorous, reproducible framework for applying BERTopic [grootendorst2022], a neural topic model, to focus group discussion transcripts. Using data from ten focus groups exploring HPV vaccine perceptions in Tunisia, we address three methodological objectives: (1) systematic exploration of hyperparameter configurations to understand their influence on model outcomes, (2) assessment of topic stability through bootstrap resampling to quantify robustness, and (3) validation of topic interpretability through human evaluation. Our work makes three primary contributions. First, we demonstrate the feasibility and value of applying transformer-based topic modeling to focus group transcripts. Second, we provide a transparent, reproducible framework for hyperparameter selection that balances multiple quality criteria including stability, coherence, and interpretability. Third, through detailed methodological reporting and comparison with traditional approaches, we contribute to the growing literature on rigorous practices in topic modeling for qualitative health research.
To ensure full reproducibility, we provide complete code, evaluation scripts, and documentation of all modeling decisions.
2 Related Work
2.1 Topic Modeling Approaches
Topic modeling encompasses a family of statistical methods for discovering latent thematic structures in document collections. Latent Dirichlet Allocation (LDA) [blei2003] established the foundational probabilistic framework for topic modeling, modeling documents as mixtures of topics and topics as distributions over words. While LDA remains widely used [jelodar2018], its bag-of-words assumption discards word order and contextual information.
BERTopic [grootendorst2022] represents a neural alternative, combining transformer-based sentence embeddings with UMAP dimensionality reduction [mcinnes2020], HDBSCAN clustering [campello2013], and class-based TF-IDF weighting. Comparative studies show BERTopic produces competitive coherence relative to LDA [egger2022], though performance depends significantly on hyperparameter choices. Recent work [borcin_optimizing_2024] demonstrated that optimal configurations vary with corpus characteristics, underscoring the importance of systematic parameter exploration.
2.2 Stability and Quality Assessment
Stability is typically quantified using partition similarity measures such as Normalized Mutual Information (NMI) or Adjusted Rand Index (ARI) [hubert1985, strehl2002]. However, these metrics measure fundamentally different aspects of clustering agreement. ARI quantifies exact pairwise assignment agreement and is sensitive to cluster boundary variations [hubert1985], while NMI measures information structure preservation based on entropy [strehl2002]. Recent systematic evaluations have demonstrated that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that neither metric is universally superior and selection should align with analytical goals. Meaney et al. [meaney2023] provide a comprehensive review of topic model quality indices and demonstrate through empirical application that different evaluation metrics often favor models of different complexity. Hosseiny and Baumer [hosseiny2023] emphasize that stability is multidimensional and that researchers must navigate tradeoffs between stability and other quality criteria.
Semantic coherence, typically measured using the $C_{v}$ metric [roder_exploring_2015], assesses whether topic words form interpretable semantic groups. However, coherence and stability represent distinct and sometimes conflicting quality dimensions [hosseiny2023, koltcov2024]. Koltcov et al. [koltcov2024] demonstrated empirically that models can exhibit high coherence while maintaining poor stability, underscoring the importance of evaluating both dimensions. Human validation remains essential for confirming that computationally optimal topics align with domain-relevant interpretations [nanda2023, steele2025].
2.3 Reproducibility and Applications to Qualitative Research
Reproducibility has emerged as a critical concern in topic modeling research. Hagg et al. [hagg2022] conducted a scoping review of 68 LDA studies in psychological science and found that most provided insufficient methodological detail to reproduce analyses, and that results are highly sensitive to analytic decisions. This highlights the need for transparent reporting of all modeling choices, systematic hyperparameter exploration, and rigorous validation.
Applications to health-related qualitative data remain relatively nascent. Studies have applied topic modeling to patient experience surveys [steele2025], psychotherapy transcripts [lalk2024], and dementia caregiver interviews [he2024], demonstrating its efficiency in identifying thematic patterns while highlighting challenges in domain-informed preprocessing and hierarchical topic organization [dou_hierarchicaltopics_2013]. Several studies have proposed augmented approaches combining computational efficiency with qualitative depth [guetterman_augmenting_2018, chang_accelerating_2021, leeson_natural_2019, crowston_using_2012].
Despite these advances, we could find no prior work explicitly applying topic modeling to focus group discussion transcripts. This gap is notable given the prevalence of focus groups in health research and the potential for topic modeling to facilitate synthesis across multiple sessions. The present study addresses this gap by developing and validating a rigorous BERTopic framework specifically designed for focus group analysis, with transparent reporting of all methodological decisions to support reproducibility and adaptation by other researchers.
3 Methodology
We apply BERTopic to HPV vaccine perception focus groups conducted in Tunisia. Our pipeline addresses three objectives: (1) systematic hyperparameter exploration, (2) stability assessment through bootstrap resampling, and (3) validation of topic interpretability. To contextualize BERTopic’s performance, we also implement a tuned Latent Dirichlet Allocation baseline for comparison (detailed in Section 3.5). Figure 1 illustrates our methodology pipeline.
Figure 1: Systematic pipeline for topic modeling of focus group transcripts. The workflow begins with transcription and translation, followed by embedding generation. A grid search explores the chosen hyperparameter configurations, evaluating each through bootstrap stability analysis (NMI, ARI) and performance metrics (coherence, topic count, outlier fraction). Multicriteria selection balances stability and interpretability, with final validation by domain experts yielding the interpretable topic structure.
3.1 Data collection and description
Between September 2023 and February 2024, we conducted ten focus group discussions exploring attitudes and perceptions toward HPV vaccination in Tunisia. The study received approval from the Institutional Review Board of Pasteur Institute of Tunis. Participants provided written informed consent, and all discussions were audio recorded and transcribed from Tunisian Dialect to French. Participants were community representatives: individuals in daily contact with children and involved in health-related decision-making, such as parents, teachers, and childcare professionals. To account for socioeconomic and cultural disparities, participants were recruited across five geographically diverse regions, with two gender-stratified sessions per region. Discussions averaged 120 minutes, were conducted in Tunisian dialect, and were moderated by a socio-anthropologist and a medical doctor using a semi-structured guide.
All discussions were audio recorded, transcribed, translated to French, and parsed into discrete rows representing individual speaker turns. Each turn was annotated with participant identifiers and enriched with metadata including gender, region, and age. Moderator questions and introductory statements were excluded to focus the analysis on substantive participant perspectives. Text preprocessing was kept minimal to preserve semantic content: we applied basic cleaning, standardized capitalization, and normalized equivalent terms (e.g., papillomavirus/HPV, coronavirus/COVID), but retained stop words to maintain contextual information for transformer-based embeddings. We removed utterances whose length was below the 5th or above the 99th percentile to remove entries that either contained too little information to be meaningfully embedded or were excessively long due to digressions or transcription artifacts.
The resulting corpus contained 1,076 speaker turns or utterances, with a mean length of 32.7 words (SD = 25.0). Table 1 presents a sample of annotated and translated speaker turns.
Original (French) Translation (English) ParticipantID Region Age
En vérité, les centres de santé de base sont toujours surpeuplés et il y a des dizaines de patients quotidiennement qui … In fact, basic health centers are always overcrowded, with dozens of patients coming in every day… par97339 Medenine 45
Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire… Excuse me, but from my personal experience and what I hear, most people have had memory problems… par39944 Bardo 47
Lorsqu’une femme souffre d’un cancer du col de l’utérus, elle préfère ne le dire à personne. Elle reste silencieuse, car… When a woman suffers from cervical cancer, she prefers not to tell anyone. She remains silent because… par73108 Sousse 43
Pourquoi ne vaccinons-nous que les filles et pas les garçons …. Why do we vaccinate only girls and not boys?… par83355 Jendouba 38
Nous sommes pour l’obligation de vaccination…. We are in favor of mandatory vaccination… par03182 Sousse 22
Je me suis disputée avec mon mari à cause du vaccin. Il travaille à l’hôpital et a essayé de me convaincre quotidienneme… I argued with my husband about the vaccine. He works at the hospital and tried to convince me every day… par49154 Medenine
Table 1: Sample of annotated and translated speaker turns
3.2 Embedding Generation and BERTopic Implementation Pipeline
We used the sentence-transformers/paraphrase-multilingual-mpnet-base-v2 model[reimers2019] to generate 768-dimensional sentence embeddings for all preprocessed sentences. This model is multilingual and optimized for sentences semantic similarity tasks, making it relevant for topic modeling in French.
Importantly, the embedding was stored and used across all subsequent model configurations, ensuring that any observed differences in topic quality or stability result from hyperparameter variations rather than inconsistencies in semantic representation. This also provides computational efficiency by eliminating redundant embedding calculations during hyperparameter exploration.
We implemented BERTopic with its standard three-component architecture: UMAP for dimensionality reduction, HDBSCAN for density-based clustering, and class-based TF-IDF for topic representation. The pipeline was configured for French language processing to ensure appropriate stop-word filtering and linguistic handling.
3.3 Hyperparameter Exploration and Stability Analysis
We conducted systematic hyperparameter exploration across three key parameters : UMAP n_neighbors, UMAP n_components, and HDBSCAN min_cluster_size. The parameter ranges were selected to span from fine-grained to coarser topic discovery: n_neighbors $\in\{10,15,30\}$, n_components $\in\{5,10,50\}$, and min_cluster_size $\in\{5,10,15\}$. To control stochasticity and ensure reproducibility, we set a global random seed that propagates through all random components of the pipeline.
For each of the 27 combinations of parameters that result, we performed the following evaluation protocol:
Bootstrap Stability Analysis:
We first generated 30 bootstrap replicas by sampling 85% of the original sentences with replacement. This sampling fraction provides sufficient data for reliable topic discovery while introducing meaningful variation to assess model robustness. Each bootstrap replicate used a deterministic random seed (base seed + replica number) to ensure reproducible sampling. For each bootstrap replicate, we instantiated fresh UMAP and HDBSCAN models. This is important because fitted dimensionality reduction and clustering models retain internal state that could bias results if reused across different data samples. Each bootstrap model was fitted independently on its respective data sample. We assessed model stability using Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), computed between all pairs of bootstrap replicates. For each metric, pairwise scores were computed only for documents appearing in both replicates, and the median value across all comparisons was retained as the stability estimate for each parameter configuration.
Full Dataset Evaluation:
We then fitted a BERTopic model on the complete dataset using the specified hyperparameters. For each fitted model, we computed topic coherence using $C_{v}$ coherence scores, recorded the number of discovered topics, and calculated the proportion of documents assigned to the outlier category.
For each parameter configuration, we recorded the hyperparameter values, bootstrap stability metrics (median NMI and median ARI), and full dataset metrics (coherence score, number of topics, outlier fraction).
3.4 Model Selection and Validation
We evaluated all parameter configurations across the above listed dimensions: stability (measured by bootstrap NMI and ARI), topic quality (coherence scores), coverage (outlier proportions), and granularity (number of topics). Rather than applying rigid filtering criteria, we examined the full parameter space to understand tradeoffs between competing objectives such as stability versus coherence, and granularity versus interpretability.
Our selection approach prioritized configurations that demonstrate strong performance across multiple criteria while producing analytically useful topic granularity. When multiple configurations showed comparable performance, we conducted detailed comparisons between them and final selection was based on coherence performance and coverage after accounting for the intended analytical requirements.
Selected configurations were validated through multiple complementary approaches. For models producing fine-grained topic structures, we performed hierarchical clustering on topic representations using cosine distance and Ward linkage. We examined the dendrogram and merged topics by selecting a cut height that produces a manageable and interpretable number of super-topics. We computed coherence scores for both original and merged topic structures to quantify the impact of hierarchical consolidation.
Finally, to assess the coherence and interpretability of the merged topics, we conducted a human validation study. Three independent annotators, with expertise in public health and qualitative analysis, rated the coherence of each topic on a 5-point Likert scale (1=Incoherent, 5=Highly Coherent) and assigned them interpretable labels. Each topic was represented with a set of randomly selected representative documents. Raters were blind to topic labels and first evaluated coherence before collaboratively assigning interpretable labels. Inter-annotator agreement was assessed using weighted Cohen’s Kappa[mchugh_interrater_2012], which accounts for the ordinal nature of the ratings, and the Intraclass Correlation Coefficient (ICC(2,1)), [shrout_intraclass_1979] which quantifies overall reliability across all three raters.
3.5 LDA Baseline Comparison
To contextualize the performance of our BERTopic approach, we implemented Latent Dirichlet Allocation as a comparative baseline. Following current practice in applied topic modeling research [hagg2022], we tuned only the number of topics while maintaining default settings for other hyperparameters. Hagg et al.’s scoping review found indeed that most applications leave parameters at defaults, making this approach representative of typical usage.
We performed grid search across the number of topics $\in\{5,7,9,10,12,15\}$, selected to span coarse to fine-grained granularity appropriate for our corpus size. Each configuration was evaluated using the same $C_{v}$ coherence scoring framework applied to BERTopic models. The optimal configuration was selected based on maximum coherence score.
4 Results
We present results from our systematic hyperparameter exploration, baseline LDA comparison, stability analysis, performance assessment, model selection, and human validation. Our evaluation spanned 27 BERTopic configurations and one LDA baseline.
4.1 Stability Analysis
We evaluated topic model stability across 27 hyperparameter combinations using bootstrap resampling with 30 replicates. Figure 2 presents stability metrics (NMI and ARI) across 27 hyperparameter combinations.
The two metrics reveal contrasting patterns across the parameter space. NMI values range from 0.376 to 0.700, with highest stability observed for min_cluster_size = 5 configurations (0.624–0.700). Stability decreases substantially for larger cluster sizes: min_cluster_size = 10 achieves 0.480–0.555, while min_cluster_size = 15 shows 0.376–0.439.
ARI values are systematically lower (0.094–0.205) and exhibit the opposite trend. Configurations with min_cluster_size = 15 achieve the highest ARI scores (0.159–0.205), while min_cluster_size = 5 shows the lowest values (0.094–0.171). This divergence reflects fundamental differences in what these metrics quantify: NMI measures information structure preservation and remains stable when fine-grained clusters consistently capture semantic distinctions, while ARI measures exact pairwise assignment agreement and is sensitive to boundary variations in small clusters where minor document shifts represent large proportional changes.
Within each min_cluster_size group, both metrics consistently show that n_neighbors = 10 produces the most stable results, while n_components has more modest effects on stability.
Figure 2: Stability metrics (NMI and ARI) across 27 hyperparameter configurations. NMI favors smaller cluster sizes while ARI shows opposite pattern, reflecting different sensitivities to cluster size.
4.2 Performance Assessment
Figure 3 shows performance metrics across the entire parameter space. Coherence scores vary substantially across configurations, ranging from 0.385 to 0.589. The highest coherence is achieved by a min_cluster_size = 15 configuration (0.589), while min_cluster_size = 5 configurations show more modest coherence values (0.385–0.427).
The number of discovered topics varies dramatically with min_cluster_size: configurations with min_cluster_size = 5 produce 36–52 topics, min_cluster_size = 10 yields 12–25 topics, and min_cluster_size = 15 generates 2–14 topics. Outlier fractions range from 5.8% to 44.1%, with generally higher outlier rates for smaller cluster size thresholds. Notably, the configuration with highest coherence (min_cluster_size = 15, n_neighbors = 30, n_components = 10) produces only 2 topics, indicating potential over-aggregation that may limit analytical utility.
Figure 3: Topic quality metrics across parameter space. Coherence scores range from 0.385-0.589, with number of topics varying from 2-52 depending on min_cluster_size.
4.3 Model Selection
Based on the stability and performance assessments, we identified two competing approaches that represent different optimization strategies:
Contender A: Stability-Optimized Approach The most stable configuration (min_cluster_size = 5, n_neighbors = 10, n_components = 5) achieved NMI = 0.695 and ARI = 0.171 and initially produced 49 topics with coherence = 0.392 and 23.6% outliers. Using hierarchical clustering analysis (Figure 4(a)), we merged related topics creating 9 semantically coherent super-topics, and improving coherence to 0.558 while maintaining 23.6% outliers.
(a) Hierarchical clustering dendrogram showing relationships among the 49 initial fine-grained topics.
(b) Composition of the 9 merged topics showing document distribution and initial fragmentation. Bar height represents total documents per topic (shown at top). White lines separate constituent subtopics, with numbers indicating document counts.
Figure 4: Hierarchical topic merging process and resulting composition. (a) The dendrogram reveals natural groupings among the 49 initial fine-grained topics extracted by BERTopic, with colored branches delineating the 9 merged topics. (b) The composition chart quantifies both the importance (bar height) and initial fragmentation (number of subtopics) of each merged topic
Contender B: Direct Coherence-Optimized Approach Among configurations producing a reasonable number of topics (>5), the highest coherence was achieved by min_cluster_size = 15, n_neighbors = 15, n_components = 10, which directly produced 9 topics with coherence = 0.539 and 26.8% outliers.
Final Selection We selected Contender A as our final model based on superior coherence performance (0.558 vs. 0.539) and lower outlier fraction (23.6% vs. 26.8%).
4.4 LDA Baseline Comparison
Systematic grid search across LDA number of topics (5-15 topics) identified 7 topics as optimal. Table 2 compares the optimized LDA baseline with our final BERTopic model.
Model Topics Coherence Outliers (%)
LDA (optimized) 7 0.546 N/A
BERTopic (merged) 9 0.558 23.6
Agreement: NMI = 0.122, ARI = 0.080
Table 2: Comparison of optimized LDA baseline and final hierarchical BERTopic model. Despite comparable coherence, minimal agreement (NMI, ARI) indicates fundamentally different topic structures.
The minimal agreement between models (NMI = 0.122, ARI = 0.080) indicates that despite comparable coherence values, LDA and BERTopic discover fundamentally different thematic structures. This divergence likely reflects architectural differences between LDA’s bag-of-words assumptions and BERTopic’s transformer-based contextual embeddings.
4.5 Final Topic Structure and Human Validation
Human validation by three domain experts confirmed that the 9 merged topics were thematically coherent and interpretable. Inter-annotator agreement was robust, with a mean pairwise weighted Cohen’s Kappa of 0.578 (moderate agreement) and a strong mean Spearman’s rank correlation ($\rho$ = 0.725). The Intraclass Correlation Coefficient (ICC) for absolute agreement was 0.79 (ICC(2,1)), indicating very good reliability per [koo_guideline_2016]. Notably, the annotators’ ratings were in perfect agreement within a one-point tolerance (100.0%), underscoring a high level of consensus on topic coherence.
Table 3 presents the coherence ratings for each topic, along with consensus labels assigned by the annotators. The mean coherence scores ranged from 2.00 to 5.00. Six of the nine topics received high ratings (mean > 4.0), including those related to Historical trust in national vaccine programs and Polarized attitudes towards official public health sources, both achieving perfect scores of 5.00. Conversely, the lowest-rated topic (mean = 2.00) concerned Prevention practices and barriers to care, suggesting that its representative sentences were perceived as less thematically unified. There was no significant correlation between a topic’s size (number of sentences) and its coherence score ($\rho$ = -0.213, p = 0.583), indicating that the perceived quality of a topic was independent of its prevalence in the discourse.
Topic ID Ann. 1 Ann. 2 Ann. 3 Mean Topic Label
Topic 0 4 4 3 3.67 COVID-19 vaccine side effects and trust erosion
Topic 1 3 4 3 3.33 Conservative norms and sexual health education
Topic 2 5 5 5 5.00 Historical trust in national vaccine programs
Topic 3 4 5 4 4.33 Parental concerns about HPV vaccination
Topic 4 5 5 4 4.67 Information sources and communication channels
Topic 5 4 4 5 4.33 Personal experiences with cancer and social impact
Topic 6 2 2 2 2.00 Prevention practices and barriers to care
Topic 7 5 5 5 5.00 Polarized attitudes towards official public health sources
Topic 8 4 5 4 4.33 General disease awareness and knowledge gaps
Inter-rater reliability: Mean weighted Cohen’s $\kappa$ = 0.578; Mean Spearman’s $\rho$ = 0.725; ICC(2,1) = 0.79
Table 3: Human validation results for the 9 merged topics. Coherence scores range from 1 (incoherent) to 5 (highly coherent). Representative utterances for each topic are provided in Appendix A.
The 9 emerging themes range from general awareness and personal experiences with cancer to specific concerns about vaccine side effects and preferred communication channels. Figure 5 visualizes the semantic organization of the topics in two-dimensional embedding space, revealing distinct thematic clusters and their spatial relationships. While this paper focuses on methodological validation, the substantive themes identified merit deeper analysis to understand their public health implications. A forthcoming companion paper will provide detailed thematic analysis and discuss policy recommendations based on these findings.
Figure 5: Two-dimensional visualization of document embeddings colored by final merged topic assignments. Each point represents a document (utterance) from the focus group transcripts. Colored regions show the 9 merged topics, with spatial proximity indicating semantic similarity.
5 Discussion
5.1 Bootstrap Stability Analysis and the NMI-ARI Divergence
Our implementation of bootstrap stability analysis builds on Greene et al. [greene2014]’s recommendations of performing bootstrap resampling for assessing topic model robustness. A critical empirical finding concerns the divergent behavior of NMI and ARI metrics across the parameter space. While NMI values ranged from 0.376 to 0.700 with highest stability at smaller cluster sizes (min_cluster_size = 5), ARI showed the opposite pattern with values of 0.094 to 0.205 and peak performance at larger cluster sizes (min_cluster_size = 15).
This divergence reflects the fundamental differences in what these metrics quantify. ARI (Adjusted Rand Index) measures exact agreement in pairwise cluster assignments [hubert1985]. It is highly sensitive to boundary shifts: when documents near cluster boundaries are reassigned across bootstrap samples, this creates pairwise disagreements that lower ARI scores. This sensitivity is particularly pronounced for small clusters, where even minor document reassignments represent large proportional changes to cluster composition. In contrast, NMI (Normalized Mutual Information) measures whether clustering solutions preserve the same underlying information structure based on entropy [strehl2002]. NMI remains relatively stable when fine-grained semantic distinctions are consistently recovered, even if exact cluster boundaries shift somewhat across bootstrap samples.
Recent systematic evaluations of topic modeling algorithms have demonstrated that ARI and NMI can favor different model configurations depending on data characteristics [rudiger_topic_2022], confirming that neither metric is universally optimal. For our exploratory analysis of focus group discussions, where the goal is to discover the full thematic landscape including both dominant and peripheral themes, we argue that NMI provides a more appropriate stability criterion. Fine-grained, information-rich models better capture the nuanced structure of focus group discourse, including smaller but important sub-themes that inform health communication interventions.
5.2 Navigating the Stability-Coherence Tradeoff
The stability-coherence tradeoff we observed, with smaller cluster sizes yielding higher stability (NMI = 0.695) but lower coherence (0.392), while larger cluster sizes produced higher coherence (0.589) but lower stability (NMI = 0.376), is an example of the multi-dimensional optimization challenge discussed by Hosseiny and Baumer [hosseiny2023], who emphasized that no single optimal configuration exists for all contexts. Our hierarchical merging strategy offers a practical solution: by extracting 49 fine-grained topics with high stability then consolidating to 9 super-topics, we achieved coherence (0.558) superior to direct extraction of 9 topics (0.539) while maintaining stability benefits.
The hierarchical approach also addresses a practical challenge: fine-grained models capture nuanced distinctions but may overwhelm readers, while coarse models sacrifice detail for simplicity. By explicitly modeling topic hierarchies, researchers can present high-level summaries while preserving access to fine-grained distinctions for detailed analysis, which is particularly valuable for focus group research where different stakeholders require different levels of granularity.
5.3 Comparison with LDA Baseline
We compared our final hierarchical BERTopic model to a systematically tuned LDA baseline. Grid search across six configurations identified 7 topics as optimal for LDA, achieving coherence of 0.5464, marginally lower than our BERTopic model (9 topics, coherence = 0.558). This small difference suggests comparable performance on semantic coherence.
However, the minimal agreement between topic assignments (NMI = 0.122, ARI = 0.080) indicates fundamentally different thematic structures. This divergence suggests that architectural differences (LDA’s bag-of-words assumptions versus BERTopic’s transformer-based embeddings) lead to qualitatively distinct topic decompositions. For conversational data where meaning depends on discourse context (e.g., "je n’ai pas confiance" referencing different trust domains), contextual embeddings may capture nuances that statistical co-occurrence patterns miss.
The near-independence of topic assignments (NMI = 0.122) suggests these methods may be complementary rather than competitive, capturing different facets of discourse structure. Future work with parallel human evaluation of both LDA and BERTopic topics could clarify whether these differences reflect genuinely distinct conceptualizations or alternative representations of similar themes.
5.4 Human Validation of Topic Quality
Our human evaluation by three independent domain experts provided essential confirmation that our computationally optimized topics align with substantive interpretations. The validation yielded robust inter-annotator agreement across multiple metrics. The Intraclass Correlation Coefficient (ICC(2,1) = 0.79) indicates very good reliability [koo_guideline_2016], demonstrating consistent evaluation across raters. Mean pairwise weighted Cohen’s kappa of 0.578 (moderate agreement) accounts for the ordinal nature of ratings while showing substantial consensus. Most strikingly, annotators achieved perfect agreement within a one-point tolerance (100%), indicating high consistency in perceived topic quality even when exact ratings differed marginally.
Six of nine topics received high ratings (mean > 4.0), with two achieving perfect scores (5.0). Only one topic received a lower rating (mean = 2.0), signaling potential for refinement. Topic size showed no correlation with coherence ($\rho$ = -0.213, p = 0.583), confirming that ratings reflected semantic quality rather than prevalence.
These findings validate that our systematic computational pipeline produces topics that domain experts recognize as interpretable and meaningful, addressing longstanding concerns [nanda2023, steele2025] that topic modeling algorithms may generate statistically optimal but substantively incoherent clusters. The strong agreement across three independent raters further suggests that the identified topics capture genuine thematic structures rather than analyst-specific interpretations. However, our validation focused exclusively on the final 9 merged topics; the 49 fine-grained initial topics were not independently evaluated. Future work could examine whether human raters prefer fine-grained versus consolidated topic structures, and whether stability-optimized configurations yield more interpretable topics than coherence-optimized alternatives.
5.5 Limitations
Several limitations merit consideration. First, the substantial sensitivity of results to hyperparameter choices raises questions about generalizability. Our finding that min_cluster_size alone shifted stability from 0.376 to 0.700, coherence from 0.385 to 0.589, and topic count from 2 to 52 demonstrates that configuration choices profoundly affect outcomes. While our systematic exploration represents a more rigorous approach than typical practice, specific optimal configurations for Tunisian HPV vaccine focus groups may not transfer to focus groups on different topics, with different conversational dynamics, or in different cultural contexts.
Second, our study is limited to French-language transcripts from Tunisia. The multilingual embedding model we employed may perform differently across languages, and the identified topic structures likely reflect culturally specific patterns of health discourse. Validation in other linguistic and cultural contexts is needed to assess the broader applicability of our methodological framework.
Third, our LDA baseline was tuned only on the number of topics following current practice [hagg2022], potentially underestimating LDA’s optimal performance. More extensive tuning of prior distributions (alpha, beta) and learning parameters could improve LDA coherence. More critically, we did not conduct parallel human evaluation of LDA topics, which would have provided direct empirical comparison of interpretability beyond algorithmic metrics.
Finally, we compared BERTopic only with LDA, not with other neural topic modeling approaches such as Top2Vec or CTM (Contextualized Topic Models). Egger and Yu [egger2022] found notable performance differences across methods depending on corpus characteristics, suggesting value in comprehensive benchmarking studies.
5.6 Implications for Focus Group Research
Despite these limitations, our work demonstrates the feasibility and value of applying rigorous topic modeling to focus group transcripts. To our knowledge, this is the first study to develop and validate a systematic topic modeling framework specifically for focus group discussions. This contribution is meaningful because focus groups remain a cornerstone methodology in qualitative health research, yet the labor-intensive nature of traditional coding limits the scale of synthesis possible. Computational approaches do not replace human interpretation but can augment it by efficiently identifying thematic patterns across multiple sessions.
Our findings suggest practical recommendations for researchers considering topic modeling for focus group analysis. First, invest in systematic hyperparameter exploration rather than accepting default configurations. Performance varies dramatically, and optimal settings depend on corpus characteristics. Second, prioritize stability assessment through bootstrap resampling. Coherence scores alone provide incomplete quality information. Third, use multiple complementary metrics (stability, coherence, coverage, granularity) and be transparent about tradeoffs in model selection. Fourth, validate computational outputs through human review by domain experts. Finally, consider hierarchical approaches that extract fine-grained topics for stability then consolidate for interpretability. This two-stage strategy can navigate the stability-coherence tradeoff effectively.
6 Conclusion
This paper presents a systematic, reproducible computational framework for applying BERTopic to focus group transcripts, contributing validated methodological practices to the computational social science toolkit. Through rigorous exploration of 27 hyperparameter configurations with bootstrap stability assessment and formal human validation, we establish that transformer-based topic modeling can reliably extract interpretable themes from structured conversational data. The strong inter-rater agreement (ICC = 0.79) confirms that computational approaches, when properly validated, produce outputs aligned with expert domain understanding.
Our work addresses a critical gap: while topic modeling has been extensively applied to social media and online text, focus group discussions, a cornerstone of qualitative research, have received limited computational attention. By demonstrating feasibility on transcripts with naturally imbalanced topic distributions and relatively small corpus size (1,076 utterances), we extend neural topic modeling to structured research settings where traditional coding remains dominant. The transparent documentation of all modeling choices, evaluation procedures, and quality tradeoffs provides a reproducible template that researchers can adapt to their own focus group contexts.
This methodological foundation enables substantive analysis of HPV vaccine perceptions in Tunisia, where our identified themes reveal specific patterns in health discourse, including vaccine hesitancy drivers, information source credibility, gender-specific concerns, and trust in health institutions. These findings will inform targeted communication strategies and policy interventions in the forthcoming companion paper. More broadly, our framework supports the integration of computational efficiency with interpretive depth, allowing researchers to scale qualitative synthesis across multiple sessions while maintaining the contextual understanding that characterizes rigorous social research. As focus groups remain central to health research, market research, and social science inquiry, validated computational methods that respect conversational data characteristics hold significant potential for advancing both methodological practice and substantive understanding.
Appendix A Appendix: Representative Topic Examples
Table 4 provides representative utterances for each of the 9 merged topics identified by our hierarchical BERTopic model, presented in both original French and English translation. These examples illustrate the thematic content and semantic coherence of each topic cluster.
Topic Label Original (French) Translation (English)
0 - COVID-19 vaccine side effects and trust erosion Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire après le vaccin contre le Covid. Je suis devenu quelqu’un qui oublie beaucoup, et d’autres confirment aussi ce constat. Moi, sincèrement, avant la crise du COVID, j’avais une confiance totale dans notre système de santé et je n’ai jamais pensé que c’était nécessaire de poser des questions, surtout que la Tunisie a une longue tradition en matière de vaccination. Mais après le COVID, je n’ai plus confiance. Le vaccin contre la grippe, je l’ai fait, c’est un peu cher, cette année je ne l’ai pas fait. Sorry, but based on my personal experience and what I hear, the majority of people have had memory problems after the Covid vaccine. I have become someone who forgets a lot, and others also confirm this observation. Honestly, before the COVID crisis, I had complete trust in our healthcare system and I never thought it was necessary to ask questions, especially since Tunisia has a long tradition of vaccination. But after COVID, I no longer have trust. I got the flu vaccine; it’s a bit expensive. This year, I didn’t get it.
1 - Conservative norms and sexual health education Vous savez, la société tunisienne est conservatrice, même lorsque quelque chose comme ça se produit, on n’en parle pas, c’est honteux. Parce que les filles sont devenues plus libres, nous avons fait une campagne de sensibilisation dans les lycées pour expliquer aux filles et aux garçons l’importance de prendre un préservatif pour se protéger. Mais, nous avons été stigmatisés. Les gens ont dit que nous sommes en train d’encourager les filles à être encore plus libres. Alors que c’est juste de la prévention. Non, en 4ème, par exemple, on lui dit que le couple est marié et que la mère est enceinte. Mais cela n’explique pas l’acte sexuel. You know, Tunisian society is conservative; even when something like that happens, people don’t talk about it, it’s shameful. Because girls have become freer, we conducted an awareness campaign in high schools to explain to girls and boys the importance of taking a condom to protect themselves. But we were stigmatized. People said that we were encouraging girls to be even freer. Whereas it is just prevention. No, in 8th grade, for example, they are told that the couple is married and that the mother is pregnant. But that does not explain the sexual act.
2 - Historical trust in national vaccine programs Les vaccins sont considérés comme un programme réussi. Lorsque vous comparez la Tunisie à des pays d’Afrique où ils sont confrontés à de nombreuses maladies graves en raison du manque de vaccins nécessaires, nous devons apprécier ce que nous avons. La vaccination en général est une bonne chose car elle protège contre la poliomyélite, l’hépatite et ceux qui ne reçoivent pas les vaccinations nécessaires peuvent avoir des problèmes de santé plus tard. Dans le public, ils ont leur carnet de vaccination. Avant de rentrer à l’école, ils doivent être vaccinés. Vaccines are considered a successful program. When you compare Tunisia to African countries where they face many serious diseases due to the lack of necessary vaccines, we must appreciate what we have. Vaccination in general is a good thing because it protects against poliomyelitis, hepatitis, and those who do not receive the necessary vaccinations may have health problems later on. In public, they have their vaccination record. Before entering school, they must be vaccinated.
3 - Parental concerns about HPV vaccination Principalement pour protéger la santé de ma fille. Comme c’est la cadette, je voulais la protéger autant que possible. Quand tu lui dis que la maladie est liée au col de l’utérus et qu’elle ne concerne que les filles, certainement que ces données le feront hésiter et penser que le vaccin pourrait peut-être affecter la fertilité de sa fille. C’est pourquoi j’insiste sur l’importance de mener des campagnes de sensibilisation pour expliquer aux gens que les filles représentent la population la plus touchée par ce virus. Mainly to protect my daughter’s health. Since she is the youngest, I wanted to protect her as much as possible. When you tell him that the disease is related to the cervix and only affects girls, these facts will certainly make him hesitate and think that the vaccine might possibly affect his daughter’s fertility. That is why I emphasize the importance of conducting awareness campaigns to explain to people that girls are the population most affected by this virus.
4 - Information sources and communication channels Les médecins doivent être formés à communiquer efficacement avec les parents. La télévision atteint un plus grand nombre de citoyens et on ne peut pas mentir à la télévision. Les gens dans les réseaux sociaux numériques publient des informations fausses et les gens finissent par les croire. Par exemple, récemment, ils ont dit qu’il y avait de la neige à Ain Draham alors qu’il n’y avait ni neige ni rien. C’est de la désinformation. Doctors must be trained to communicate effectively with parents. Television reaches a larger number of citizens and one cannot lie on television. People on digital social networks publish false information and people end up believing it. For example, recently, they said there was snow in Ain Draham when there was neither snow nor anything. This is misinformation.
5 - Personal experiences with cancer and its social impact Je connais quelqu’un qui en a souffert et qui a tout enlevé, elle avait 38 ans, elle en a été affectée et déprimée. Elle est mariée et a un enfant, c’est ma cousine, elle a été hospitalisée, elle a subi une opération pour enlever son utérus. Il y a aussi une maladie très répandue qui est le cancer du foie. J’ai subi deux opérations chirurgicales pour l’enlever. N’y a-t-il pas de vaccination contre cela ? J’ai une voisine qui est tombée malade de cette maladie et son mari l’a divorcée. Les hommes ne soutiennent pas les femmes dans ces situations. I know someone who suffered from it and had everything removed; she was 38 years old, and it affected her deeply and made her depressed. She is married and has a child; she is my cousin. She was hospitalized and underwent surgery to have her uterus removed. There is also a very common disease which is liver cancer. I have undergone two surgical operations to remove it. Is there no vaccination against it? I have a neighbor who fell ill with this disease and her husband divorced her. Men do not support women in these situations.
6 - Prevention practices and barriers to care Moi, je ne veux pas faire le frottis car j’ai peur. J’ai eu une mauvaise expérience après mon accouchement dans un hôpital public et je ne veux pas que quiconque me touche dans ma zone intime. Oui, je consulte mon médecin chaque année pour un contrôle. Il y a des remèdes naturels que j’utilise quand mon enfant est malade, comme le miel et l’eucalyptus. I don’t want to have the smear test because I’m afraid. I had a bad experience after my delivery in a public hospital and I don’t want anyone to touch me in my intimate area. Yes, I see my doctor every year for a check-up. There are natural remedies that I use when my child is sick, such as honey and eucalyptus.
7 - Polarized attitudes towards official public health sources Je ne sais pas, je n’ai pas confiance dans l’ensemble du système. Nous faisons confiance aux vaccins approuvés par l’État et le ministère de la Santé. Moi, je fais confiance à la page officielle du ministère de la Santé. I don’t know, I don’t trust the whole system. We trust vaccines approved by the state and the Ministry of Health. I trust the official page of the Ministry of Health.
8 - General disease awareness and knowledge gaps C’est la première fois que j’entends parler de cette maladie. L’hygiène est l’une des principales causes de cette maladie. Je n’ai entendu parler d’aucun cas, je sais seulement que c’est une maladie maligne. This is the first time I have heard of this disease. Hygiene is one of the main causes of this disease. I have not heard of any cases; I only know that it is a malignant disease.
Table 4: Representative utterances for each of the 9 merged topics with original French text and English translations.

### v2

@@
 Focus group discussions generate rich qualitative data but their analysis traditionally relies on labor-intensive manual coding that limits scalability and reproducibility.
-We present a rigorous, reproducible computational framework for applying neural topic modeling to focus group transcripts, addressing fundamental methodological challenges: hyperparameter sensitivity, model stability, and validation of interpretability.
-Using BERTopic applied to ten focus groups exploring HPV vaccine perceptions in Tunisia (1,076 utterances), we conducted systematic evaluation across 27 hyperparameter configurations, assessed stability through bootstrap resampling with 30 replicates per configuration, and validated interpretability through formal human evaluation by three domain experts.
-Our analysis demonstrates substantial sensitivity to hyperparameter choices and reveals that metric selection for stability assessment must align with analytical goals.
-A hierarchical merging strategy (extracting fine-grained topics for stability then consolidating for interpretability) effectively navigates the stability-coherence tradeoff, achieving coherence of 0.558 compared to 0.539 for direct extraction.
-Human validation confirmed topic quality with very good inter-rater reliability (ICC = 0.79, weighted Cohen’s kappa = 0.578).
-Our framework provides practical guidelines that researchers can adapt to their own qualitative research contexts.
-All code, data processing scripts, and evaluation protocols are publicly available to support reproduction and extension of this work.
+We present a systematic framework for applying BERTopic to focus group transcripts using data from ten focus groups exploring HPV vaccine perceptions in Tunisia (1,075 utterances).
+We conducted comprehensive hyperparameter exploration across 27 configurations, evaluating each through bootstrap stability analysis, performance metrics, and comparison with LDA baseline.
+Bootstrap analysis revealed that stability metrics (NMI and ARI) exhibited strong disagreement (r = -0.691) and showed divergent relationships with coherence, demonstrating that stability is multifaceted rather than monolithic.
+Our multi-criteria selection framework yielded a 7-topic model achieving 18% higher coherence than optimized LDA (0.573 vs.
+0.486) with interpretable topics validated through independent human evaluation (ICC = 0.700, weighted Cohen’s $\kappa$ = 0.678).
+These findings demonstrate that transformer-based topic modeling can extract interpretable themes from small focus group transcript corpora when systematically configured and validated, while revealing that quality metrics capture distinct, sometimes conflicting constructs requiring multi-criteria evaluation.
+We provide complete documentation and code to support reproducibility.
 Keywords Neural Topic Modeling $\cdot$ BERTopic $\cdot$ Focus Group Analysis $\cdot$ Model Validation $\cdot$ Reproducibility $\cdot$ Hyperparameter Optimization $\cdot$ Bootstrap Resampling $\cdot$ Text Mining
@@
 Topic modeling has emerged as a powerful computational approach for discovering latent thematic structures in large text corpora.
-Originally developed for document organization and information retrieval, topic modeling techniques have found increasing application in public health research, where understanding discourse patterns can inform intervention strategies and policy decisions.
-The COVID-19 pandemic accelerated this trend, with numerous studies applying topic modeling to analyze public sentiment and concerns expressed on social media platforms [qorib_covid-19_2023, figueiredo_topic_2024, yin_sentiment_2022, lyu_covid-19_2021, boon-itt_public_2020].
-Similarly, topic modeling has been deployed to characterize discussions about Human Papillomaviruss (HPV) vaccination [du_use_2020, surian_characterizing_2016, rai_detecting_2024], cancer care experiences [steele2025], and mental health support communities [carron-arthur_whats_2016].
-These applications demonstrate the growing recognition of topic modeling as a valuable tool for public health surveillance and understanding health-related discourse at scale.
-However, the majority of these applications focus on readily available online data sources such as Twitter, Reddit, and online forums.
-While social media provides unprecedented access to large-scale public discourse, it represents only one facet of health communication.
-Focus group discussions, a cornerstone methodology in qualitative health research, offer structured settings where participants engage in moderated conversations that elicit deeper, contextualized perspectives on health beliefs, barriers, and decision-making processes.
-Despite their methodological importance, focus group transcripts have received limited attention in the topic modeling literature.
-This gap likely reflects both the proprietary nature of focus group data and the traditional emphasis on human-coded qualitative analysis.
-Yet focus groups generate rich textual data that could benefit from computational approaches, particularly when research objectives include systematic synthesis across multiple sessions or geographic contexts.
-The application of topic modeling to qualitative research raises important methodological challenges.
-Focus group transcripts are typically small, heterogeneous corpora with non-uniform conversational structure, which can push topic models outside of the conditions under which they are commonly benchmarked.
-In such low-data regimes, model outcomes can be highly sensitive to design choices.
-At the same time, topic modeling algorithms are inherently stochastic, which means that seemingly minor configuration changes can lead to very different topic structures which raises questions about the reproducibility and stability of results, especially when findings inform public health interventions.
-Recent methodological reviews have highlighted widespread issues with reproducibility in topic modeling research [hagg2022] and emphasized the multidimensional nature of model quality [hosseiny2023].
-This paper presents a rigorous, reproducible framework for applying BERTopic [grootendorst2022], a neural topic model, to focus group discussion transcripts.
-Using data from ten focus groups exploring HPV vaccine perceptions in Tunisia, we address three methodological objectives: (1) systematic exploration of hyperparameter configurations to understand their influence on model outcomes, (2) assessment of topic stability through bootstrap resampling to quantify robustness, and (3) validation of topic interpretability through human evaluation.
-Our work makes three primary contributions.
-First, we demonstrate the feasibility and value of applying transformer-based topic modeling to focus group transcripts.
-Second, we provide a transparent, reproducible framework for hyperparameter selection that balances multiple quality criteria including stability, coherence, and interpretability.
-Third, through detailed methodological reporting and comparison with traditional approaches, we contribute to the growing literature on rigorous practices in topic modeling for qualitative health research.
-To ensure full reproducibility, we provide complete code, evaluation scripts, and documentation of all modeling decisions.
+Originally developed for document organization and information retrieval [blei2003], topic modeling techniques have found widespread application across diverse text types.
+Early work focused on structured document collections such as news articles, scientific papers, and web pages [griffiths2004, newman2006].
+More recently, researchers have applied topic modeling extensively to social media platforms including Twitter [qorib_covid-19_2023, lyu_covid-19_2021], Reddit [fong2025, yang2025], and online forums [sik2023, davila2023], as well as to customer reviews [yazici2024, liu2023] and open-ended survey responses [guetterman_augmenting_2018].
+Recent advances in transformer-based architectures have enabled neural topic models that leverage contextual embeddings rather than traditional bag-of-words representations, offering improved performance for short, noisy text common in social media [grootendorst2022, egger2022].
+Despite this breadth of applications, most implementations target large-scale corpora where statistical inference benefits from thousands or tens of thousands of documents.
+Moreover, standard benchmarks and evaluation protocols emphasize either well-structured documents (e.g., 20 Newsgroups, scientific abstracts) or homogeneous short texts (e.g., tweets about specific events).
+This leaves open questions about topic modeling performance on small, heterogeneous conversational corpora where documents vary substantially in length and thematic focus, and where corpus size may comprise only hundreds of utterances.
+Focus group discussions represent a particularly important yet computationally underexplored text type.
+Unlike social media posts, which are spontaneous and often brief, focus groups are structured conversations where participants engage in moderated dialogue following a semi-structured protocol.
+Unlike customer reviews or survey responses, which target specific products or questions, focus groups explore open-ended topics through emergent conversation.
+Focus groups are widely used across health research, market research, policy analysis, and user experience studies, generating rich textual data that captures nuanced perspectives through natural conversation.
+Despite their methodological prevalence, focus group transcripts have received limited attention in the computational topic modeling literature, likely reflecting both the proprietary nature of such data and the traditional emphasis on manual thematic coding.
+Yet when research objectives include systematic synthesis across multiple sessions, geographic contexts, or longitudinal studies, computational approaches could augment traditional analysis by efficiently identifying thematic patterns while preserving opportunities for human interpretation.
+This paper presents a systematic, reproducible framework for applying BERTopic [grootendorst2022], a neural topic model, to focus group discussion transcripts.
+Using data from ten focus groups exploring HPV vaccine perceptions in Tunisia, we address four methodological objectives: (1) systematic exploration of hyperparameter configurations to understand their influence on model outcomes, (2) assessment of topic stability through bootstrap resampling to quantify robustness, (3) comparison with Latent Dirichlet Allocation to contextualize BERTopic’s performance against traditional approaches, and (4) validation of topic interpretability through independent human evaluation by domain experts.
+While our application focuses on health discourse, the methodological framework and findings have broader implications for topic modeling of small, heterogeneous conversational corpora common across qualitative research domains.
+Our work makes three primary contributions to computational text analysis methodology.
+First, we develop and validate the first systematic framework specifically designed for topic modeling of focus group transcripts, demonstrating that transformer-based approaches can extract interpretable themes from conversational qualitative data when properly configured and validated.
+Second, we show that BERTopic substantially outperforms LDA for this data type, achieving 18% higher coherence while discovering more interpretable topics, with human validation confirming that computational coherence metrics successfully predict expert judgments.
+Third, through transparent documentation of all modeling decisions, systematic evaluation across multiple quality dimensions, and rigorous comparison with traditional approaches, we contribute to the growing literature on reproducible practices in qualitative research.
+To ensure full reproducibility, we provide complete code, evaluation scripts, hyperparameter configurations, and detailed documentation of all modeling decisions.
+Code and anonymized data are available at ….
 2 Related Work
@@
 2.2 Stability and Quality Assessment
+Assessing topic model quality requires evaluating multiple dimensions, as different metrics capture distinct aspects of model performance.
+Greene et al.
+[greene2014] advocated for bootstrap stability analysis as a criterion for topic model selection, proposing that models producing stable partitions across resampled data represent more reliable thematic structures.
 Stability is typically quantified using partition similarity measures such as Normalized Mutual Information (NMI) or Adjusted Rand Index (ARI) [hubert1985, strehl2002].
@@
 Recent systematic evaluations have demonstrated that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that neither metric is universally superior and selection should align with analytical goals.
-Meaney et al.
-[meaney2023] provide a comprehensive review of topic model quality indices and demonstrate through empirical application that different evaluation metrics often favor models of different complexity.
-Hosseiny and Baumer [hosseiny2023] emphasize that stability is multidimensional and that researchers must navigate tradeoffs between stability and other quality criteria.
 Semantic coherence, typically measured using the $C_{v}$ metric [roder_exploring_2015], assesses whether topic words form interpretable semantic groups.
@@
 [koltcov2024] demonstrated empirically that models can exhibit high coherence while maintaining poor stability, underscoring the importance of evaluating both dimensions.
-Human validation remains essential for confirming that computationally optimal topics align with domain-relevant interpretations [nanda2023, steele2025].
-2.3 Reproducibility and Applications to Qualitative Research
-Reproducibility has emerged as a critical concern in topic modeling research.
+Furthermore, algorithmic metrics frequently conflict with human judgments of topic quality [hoyle2021, chang2009], and different coherence measures can disagree on model rankings [aletras2013].
+Meaney et al.
+[meaney2023] demonstrate that different evaluation metrics often favor models of different complexity, indicating that no single metric captures all aspects of topic quality.
+These findings underscore the importance of multi-faceted evaluation that includes expert validation alongside computational metrics.
+2.3 Applications to Qualitative Research and Reproducibility Challenges
+Topic modeling has found increasing application in qualitative research across multiple domains.
+Substantial work has applied these methods to social media health discourse, including COVID-19 discussions on Twitter [qorib_covid-19_2023] and HPV vaccination debates [du_use_2020].
+More recently, researchers have analyzed structured qualitative data including patient experience surveys [steele2025], psychotherapy transcripts [lalk2024], interview data [he2024], and open-ended survey responses [guetterman_augmenting_2018, chang_accelerating_2021].
+These applications demonstrate topic modeling’s potential for efficiently identifying thematic patterns in qualitative data while highlighting methodological challenges including domain-informed preprocessing, handling conversational dynamics, and hierarchical topic organization [dou_hierarchicaltopics_2013].
+However, reproducibility has emerged as a critical concern.
 Hagg et al.
 [hagg2022] conducted a scoping review of 68 LDA studies in psychological science and found that most provided insufficient methodological detail to reproduce analyses, and that results are highly sensitive to analytic decisions.
-This highlights the need for transparent reporting of all modeling choices, systematic hyperparameter exploration, and rigorous validation.
-Applications to health-related qualitative data remain relatively nascent.
-Studies have applied topic modeling to patient experience surveys [steele2025], psychotherapy transcripts [lalk2024], and dementia caregiver interviews [he2024], demonstrating its efficiency in identifying thematic patterns while highlighting challenges in domain-informed preprocessing and hierarchical topic organization [dou_hierarchicaltopics_2013].
-Several studies have proposed augmented approaches combining computational efficiency with qualitative depth [guetterman_augmenting_2018, chang_accelerating_2021, leeson_natural_2019, crowston_using_2012].
-Despite these advances, we could find no prior work explicitly applying topic modeling to focus group discussion transcripts.
-This gap is notable given the prevalence of focus groups in health research and the potential for topic modeling to facilitate synthesis across multiple sessions.
-The present study addresses this gap by developing and validating a rigorous BERTopic framework specifically designed for focus group analysis, with transparent reporting of all methodological decisions to support reproducibility and adaptation by other researchers.
+This challenge is particularly acute for small, heterogeneous qualitative corpora where configuration choices can profoundly affect outcomes.
+The lack of standardized evaluation protocols and transparent reporting practices undermines confidence in topic modeling findings, especially when results inform practical decisions.
+These issues highlight the need for systematic hyperparameter exploration, multi-dimensional quality assessment, rigorous validation protocols, and transparent documentation of all modeling choices.
+Despite growing applications, we could find no prior work explicitly applying topic modeling to focus group discussion transcripts.
+This gap is notable: focus groups generate multi-party conversations where meaning emerges through interactive dialogue, differing substantively from both social media (spontaneous, brief posts) and individual interviews (single-voice narratives).
+The present study addresses this gap by developing and validating a systematic BERTopic framework for focus group analysis, with transparent reporting of all methodological decisions to support reproducibility and adaptation by other researchers.
 3 Methodology
@@
 Our pipeline addresses three objectives: (1) systematic hyperparameter exploration, (2) stability assessment through bootstrap resampling, and (3) validation of topic interpretability.
-To contextualize BERTopic’s performance, we also implement a tuned Latent Dirichlet Allocation baseline for comparison (detailed in Section 3.5).
+To contextualize BERTopic’s performance, we also implement a tuned LDA baseline for comparison (detailed in Section 3.6).
 Figure 1 illustrates our methodology pipeline.
@@
 A grid search explores the chosen hyperparameter configurations, evaluating each through bootstrap stability analysis (NMI, ARI) and performance metrics (coherence, topic count, outlier fraction).
-Multicriteria selection balances stability and interpretability, with final validation by domain experts yielding the interpretable topic structure.
+Multicriteria selection balances performance metrics, with final validation by domain experts giving the interpretable topic structure.
 3.1 Data collection and description
@@
 We removed utterances whose length was below the 5th or above the 99th percentile to remove entries that either contained too little information to be meaningfully embedded or were excessively long due to digressions or transcription artifacts.
-The resulting corpus contained 1,076 speaker turns or utterances, with a mean length of 32.7 words (SD = 25.0). Table 1 presents a sample of annotated and translated speaker turns.
+The resulting corpus contained 1,075 speaker turns or utterances, with a mean length of 32.7 words (SD = 25.0). Table 1 presents a sample of annotated and translated speaker turns.
 Original (French) Translation (English) ParticipantID Region Age
-En vérité, les centres de santé de base sont toujours surpeuplés et il y a des dizaines de patients quotidiennement qui … In fact, basic health centers are always overcrowded, with dozens of patients coming in every day… par97339 Medenine 45
-Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire… Excuse me, but from my personal experience and what I hear, most people have had memory problems… par39944 Bardo 47
+En vérité, les centres de santé de base sont toujours surpeuplés et il y a des dizaines de patients quotidiennement qui … In fact, basic health centers are always overcrowded, with dozens of patients coming in every day… par97339 Region1 45
+Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire… Excuse me, but from my personal experience and what I hear, most people have had memory problems… par39944 Region2 47
 Lorsqu’une femme souffre d’un cancer du col de l’utérus, elle préfère ne le dire à personne.
 Elle reste silencieuse, car… When a woman suffers from cervical cancer, she prefers not to tell anyone.
-She remains silent because… par73108 Sousse 43
-Pourquoi ne vaccinons-nous que les filles et pas les garçons …. Why do we vaccinate only girls and not boys?… par83355 Jendouba 38
-Nous sommes pour l’obligation de vaccination…. We are in favor of mandatory vaccination… par03182 Sousse 22
+She remains silent because… par73108 Region3 43
+Pourquoi ne vaccinons-nous que les filles et pas les garçons …. Why do we vaccinate only girls and not boys?… par83355 Region4 38
+Nous sommes pour l’obligation de vaccination…. We are in favor of mandatory vaccination… par03182 Region3 22
 Je me suis disputée avec mon mari à cause du vaccin.
 Il travaille à l’hôpital et a essayé de me convaincre quotidienneme… I argued with my husband about the vaccine.
-He works at the hospital and tried to convince me every day… par49154 Medenine
-Table 1: Sample of annotated and translated speaker turns
+He works at the hospital and tried to convince me every day… par49154 Region1
+Table 1: Sample of annotated and translated speaker turns (utterances). Participants names and regions have been anonymized and the dataset doesn’t contain any Personally Identifiable Information
 3.2 Embedding Generation and BERTopic Implementation Pipeline
 We used the sentence-transformers/paraphrase-multilingual-mpnet-base-v2 model[reimers2019] to generate 768-dimensional sentence embeddings for all preprocessed sentences.
-This model is multilingual and optimized for sentences semantic similarity tasks, making it relevant for topic modeling in French.
-Importantly, the embedding was stored and used across all subsequent model configurations, ensuring that any observed differences in topic quality or stability result from hyperparameter variations rather than inconsistencies in semantic representation.
+This model is multilingual and optimized for sentence semantic similarity tasks, making it relevant for topic modeling in French.
+Importantly, the embeddings were computed once and stored for use across all subsequent model configurations, ensuring that any observed differences in topic quality or stability result from hyperparameter variations rather than inconsistencies in semantic representation.
 This also provides computational efficiency by eliminating redundant embedding calculations during hyperparameter exploration.
@@
 To control stochasticity and ensure reproducibility, we set a global random seed that propagates through all random components of the pipeline.
-For each of the 27 combinations of parameters that result, we performed the following evaluation protocol:
-Bootstrap Stability Analysis:
-We first generated 30 bootstrap replicas by sampling 85% of the original sentences with replacement.
+For each of the 27 resulting parameter combinations, we performed the following evaluation protocol:
+Bootstrap Stability Analysis: We generated 30 bootstrap replicates by sampling 85% of the original sentences with replacement.
 This sampling fraction provides sufficient data for reliable topic discovery while introducing meaningful variation to assess model robustness.
-Each bootstrap replicate used a deterministic random seed (base seed + replica number) to ensure reproducible sampling.
+Each bootstrap replicate used a deterministic random seed (base seed + replicate number) to ensure reproducible sampling.
 For each bootstrap replicate, we instantiated fresh UMAP and HDBSCAN models.
-This is important because fitted dimensionality reduction and clustering models retain internal state that could bias results if reused across different data samples.
+This is essential because fitted dimensionality reduction and clustering models retain internal state that could bias results if reused across different data samples.
 Each bootstrap model was fitted independently on its respective data sample.
 We assessed model stability using Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), computed between all pairs of bootstrap replicates.
-For each metric, pairwise scores were computed only for documents appearing in both replicates, and the median value across all comparisons was retained as the stability estimate for each parameter configuration.
-Full Dataset Evaluation:
-We then fitted a BERTopic model on the complete dataset using the specified hyperparameters.
+For each metric, pairwise scores were computed only for documents appearing in both replicates, and the median value across all pairwise comparisons was retained as the stability estimate for that parameter configuration.
+Full Dataset Evaluation: We fitted a BERTopic model on the complete dataset using each parameter configuration.
 For each fitted model, we computed topic coherence using $C_{v}$ coherence scores, recorded the number of discovered topics, and calculated the proportion of documents assigned to the outlier category.
-For each parameter configuration, we recorded the hyperparameter values, bootstrap stability metrics (median NMI and median ARI), and full dataset metrics (coherence score, number of topics, outlier fraction).
-3.4 Model Selection and Validation
-We evaluated all parameter configurations across the above listed dimensions: stability (measured by bootstrap NMI and ARI), topic quality (coherence scores), coverage (outlier proportions), and granularity (number of topics).
-Rather than applying rigid filtering criteria, we examined the full parameter space to understand tradeoffs between competing objectives such as stability versus coherence, and granularity versus interpretability.
-Our selection approach prioritized configurations that demonstrate strong performance across multiple criteria while producing analytically useful topic granularity.
-When multiple configurations showed comparable performance, we conducted detailed comparisons between them and final selection was based on coherence performance and coverage after accounting for the intended analytical requirements.
-Selected configurations were validated through multiple complementary approaches.
-For models producing fine-grained topic structures, we performed hierarchical clustering on topic representations using cosine distance and Ward linkage.
-We examined the dendrogram and merged topics by selecting a cut height that produces a manageable and interpretable number of super-topics.
-We computed coherence scores for both original and merged topic structures to quantify the impact of hierarchical consolidation.
-Finally, to assess the coherence and interpretability of the merged topics, we conducted a human validation study.
-Three independent annotators, with expertise in public health and qualitative analysis, rated the coherence of each topic on a 5-point Likert scale (1=Incoherent, 5=Highly Coherent) and assigned them interpretable labels.
-Each topic was represented with a set of randomly selected representative documents.
-Raters were blind to topic labels and first evaluated coherence before collaboratively assigning interpretable labels.
-Inter-annotator agreement was assessed using weighted Cohen’s Kappa[mchugh_interrater_2012], which accounts for the ordinal nature of the ratings, and the Intraclass Correlation Coefficient (ICC(2,1)), [shrout_intraclass_1979] which quantifies overall reliability across all three raters.
+3.4 Multi-Criteria Model Selection
+Rather than applying rigid optimization criteria or automated selection rules, we adopted a multi-criteria decision framework balancing four dimensions:
+  • Coherence: Semantic interpretability of topics, measured by $C_{v}$ scores
+  • Stability: Bootstrap consistency, assessed through NMI and ARI
+  • Granularity: Number of non-outlier topics, determining thematic resolution
+  • Coverage: Proportion of documents assigned to topics versus outliers
+This approach acknowledges that optimal configurations depend on research context: exploratory qualitative research may prioritize coherence, while studies requiring reproducible classifications emphasize stability.
+By transparently presenting performance across all 27 configurations, we enable informed selection rather than imposing a single optimization objective.
+3.5 Human Validation
+To assess the coherence and interpretability of our final topic model, we conducted a validation study with three independent annotators having expertise in public health and qualitative analysis.
+Each annotator rated the coherence of each topic on a 5-point Likert scale (1=Incoherent, 5=Highly Coherent) based on 10 representative documents randomly selected from each topic.
+Raters were blind to topic labels and model-generated keywords, and first evaluated coherence independently before collaboratively assigning interpretable labels to each topic.
+Inter-annotator agreement was assessed using weighted Cohen’s Kappa[mchugh_interrater_2012], which accounts for the ordinal nature of ratings, and the Intraclass Correlation Coefficient (ICC(2,1))[shrout_intraclass_1979], which quantifies overall reliability across all three raters.
 3.6 LDA Baseline Comparison
@@
 Following current practice in applied topic modeling research [hagg2022], we tuned only the number of topics while maintaining default settings for other hyperparameters.
-Hagg et al.’s scoping review found indeed that most applications leave parameters at defaults, making this approach representative of typical usage.
-We performed grid search across the number of topics $\in\{5,7,9,10,12,15\}$, selected to span coarse to fine-grained granularity appropriate for our corpus size.
+Hagg et al.’s scoping review found that most applications leave parameters at defaults, making this approach representative of typical usage.
+We performed grid search across the number of topics $\in\{3,\cdots,15\}$, selected to span coarse to fine-grained granularity appropriate for our corpus size.
 Each configuration was evaluated using the same $C_{v}$ coherence scoring framework applied to BERTopic models.
-The optimal configuration was selected based on maximum coherence score.
+The optimal LDA configuration was selected based on maximum coherence score.
 4 Results
-We present results from our systematic hyperparameter exploration, baseline LDA comparison, stability analysis, performance assessment, model selection, and human validation.
-Our evaluation spanned 27 BERTopic configurations and one LDA baseline.
 4.1 Stability Analysis
-We evaluated topic model stability across 27 hyperparameter combinations using bootstrap resampling with 30 replicates.
-Figure 2 presents stability metrics (NMI and ARI) across 27 hyperparameter combinations.
-The two metrics reveal contrasting patterns across the parameter space.
-NMI values range from 0.376 to 0.700, with highest stability observed for min_cluster_size = 5 configurations (0.624–0.700).
-Stability decreases substantially for larger cluster sizes: min_cluster_size = 10 achieves 0.480–0.555, while min_cluster_size = 15 shows 0.376–0.439.
-ARI values are systematically lower (0.094–0.205) and exhibit the opposite trend.
-Configurations with min_cluster_size = 15 achieve the highest ARI scores (0.159–0.205), while min_cluster_size = 5 shows the lowest values (0.094–0.171).
-This divergence reflects fundamental differences in what these metrics quantify: NMI measures information structure preservation and remains stable when fine-grained clusters consistently capture semantic distinctions, while ARI measures exact pairwise assignment agreement and is sensitive to boundary variations in small clusters where minor document shifts represent large proportional changes.
-Within each min_cluster_size group, both metrics consistently show that n_neighbors = 10 produces the most stable results, while n_components has more modest effects on stability.
-Figure 2: Stability metrics (NMI and ARI) across 27 hyperparameter configurations.
-NMI favors smaller cluster sizes while ARI shows opposite pattern, reflecting different sensitivities to cluster size.
+We evaluated topic model stability across 27 hyperparameter combinations using bootstrap resampling with 30 replicates per configuration.
+This analysis showed NMI values ranging from 0.388 to 0.703 (median: 0.558), while ARI values ranged from 0.107 to 0.202 (median: 0.156).
+Figure 2 (top panels) presents these stability metrics across the parameter space.
+The two metrics reveal contrasting patterns.
+NMI values are highest for min_cluster_size = 5 configurations (0.63-0.70), decreasing substantially for larger cluster sizes: min_cluster_size = 10 achieves 0.48-0.56, while min_cluster_size = 15 shows 0.39-0.44.
+ARI values are systematically lower and exhibit the opposite trend: configurations with min_cluster_size = 15 achieve the highest ARI scores (0.15-0.20), while min_cluster_size = 5 shows the lowest values (0.11-0.16).
+This divergence reflects fundamental differences in what these metrics quantify.
+NMI measures information structure preservation and remains stable when fine-grained clusters consistently capture semantic distinctions, even when exact boundaries shift across bootstrap samples.
+ARI measures exact pairwise assignment agreement and is highly sensitive to boundary variations, particularly in small clusters where minor document shifts represent large proportional changes.
+The negative correlation between NMI and ARI (r = -0.691) further illustrates their divergent behavior.
+Critically, we observed a strong negative correlation between NMI and model coherence (r = -0.911, p < 0.001).
+Configurations with highest NMI (0.63-0.70) produced the least coherent topics (coherence = 0.39-0.43), while configurations with lower NMI (0.39-0.44) achieved substantially higher coherence (0.52-0.59).
+In contrast, ARI demonstrated a moderate positive correlation with coherence (r = 0.639, p < 0.001), though the universally low ARI range limits its practical utility for model selection.
+These patterns indicate that stability as measured by NMI and semantic quality as measured by coherence capture fundamentally different constructs, with implications for model selection strategies.
 4.2 Performance Assessment
-Figure 3 shows performance metrics across the entire parameter space.
-Coherence scores vary substantially across configurations, ranging from 0.385 to 0.589.
-The highest coherence is achieved by a min_cluster_size = 15 configuration (0.589), while min_cluster_size = 5 configurations show more modest coherence values (0.385–0.427).
-The number of discovered topics varies dramatically with min_cluster_size: configurations with min_cluster_size = 5 produce 36–52 topics, min_cluster_size = 10 yields 12–25 topics, and min_cluster_size = 15 generates 2–14 topics.
-Outlier fractions range from 5.8% to 44.1%, with generally higher outlier rates for smaller cluster size thresholds.
-Notably, the configuration with highest coherence (min_cluster_size = 15, n_neighbors = 30, n_components = 10) produces only 2 topics, indicating potential over-aggregation that may limit analytical utility.
-Figure 3: Topic quality metrics across parameter space. Coherence scores range from 0.385-0.589, with number of topics varying from 2-52 depending on min_cluster_size.
+Figure 2 (middle and bottom panels) displays performance metrics across the entire parameter space.
+Coherence scores vary substantially, ranging from 0.387 to 0.587 (mean = 0.478, SD = 0.060).
+The highest coherence is achieved by coarser configurations with min_cluster_size = 15 (0.50-0.59), while min_cluster_size = 5 configurations show lower coherence (0.39-0.43) despite higher stability by NMI.
+Figure 2: Quality metrics across the complete hyperparameter space.
+Six metrics are displayed as heatmaps across three minimum cluster sizes (columns).
+Each heatmap shows UMAP n_neighbors (rows: 10, 15, 30) by UMAP n_components (columns: 5, 10, 50).
+Top panels show bootstrap stability (NMI and ARI), middle panels show topic quality (coherence and number of topics), and bottom panels show coverage metrics (outlier fraction and main topic proportion).
+The number of discovered topics varies dramatically with min_cluster_size: configurations with min_cluster_size = 5 produce 34-55 topics, min_cluster_size = 10 yields 10-25 topics, and min_cluster_size = 15 generates 2-12 topics.
+Outlier fractions range from 0.7% to 38.8%, with generally higher outlier rates for finer-grained configurations.
+Main topic proportions range from 4% to 66%, indicating substantial variation in topic size distributions.
 4.3 Model Selection
-Based on the stability and performance assessments, we identified two competing approaches that represent different optimization strategies:
-Contender A: Stability-Optimized Approach The most stable configuration (min_cluster_size = 5, n_neighbors = 10, n_components = 5) achieved NMI = 0.695 and ARI = 0.171 and initially produced 49 topics with coherence = 0.392 and 23.6% outliers.
-Using hierarchical clustering analysis (Figure 4(a)), we merged related topics creating 9 semantically coherent super-topics, and improving coherence to 0.558 while maintaining 23.6% outliers.
-(a) Hierarchical clustering dendrogram showing relationships among the 49 initial fine-grained topics.
-(b) Composition of the 9 merged topics showing document distribution and initial fragmentation.
-Bar height represents total documents per topic (shown at top).
-White lines separate constituent subtopics, with numbers indicating document counts.
-Figure 4: Hierarchical topic merging process and resulting composition.
-(a) The dendrogram reveals natural groupings among the 49 initial fine-grained topics extracted by BERTopic, with colored branches delineating the 9 merged topics.
-(b) The composition chart quantifies both the importance (bar height) and initial fragmentation (number of subtopics) of each merged topic
-Contender B: Direct Coherence-Optimized Approach Among configurations producing a reasonable number of topics (>5), the highest coherence was achieved by min_cluster_size = 15, n_neighbors = 15, n_components = 10, which directly produced 9 topics with coherence = 0.539 and 26.8% outliers.
-Final Selection We selected Contender A as our final model based on superior coherence performance (0.558 vs. 0.539) and lower outlier fraction (23.6% vs. 26.8%).
+Table 2 presents the top 10 configurations ranked by coherence score.
+Rank $n$ nc mcs Topics Coherence NMI ARI Outliers
+1 10 10 15 2 0.587 0.435 0.167 0.007
+2* 15 10 15 7 0.573 0.388 0.164 0.233
+3 10 50 15 8 0.549 0.420 0.165 0.249
+4 30 10 15 10 0.523 0.420 0.195 0.356
+5 30 5 10 14 0.521 0.507 0.146 0.388
+6 15 10 10 11 0.519 0.531 0.151 0.286
+7 30 50 15 10 0.516 0.419 0.189 0.371
+8 15 5 15 10 0.515 0.416 0.153 0.236
+9 30 5 15 9 0.511 0.420 0.202 0.335
+10 10 5 15 12 0.504 0.431 0.163 0.367
+Table 2: Top 10 configurations ranked by coherence score.
+Configurations are identified by base hyperparameters (n_neighbors ($n$), n_components(nc), and min_cluster_size (mcs)).
+The selected configuration (rank 2, marked with *) balances high coherence with reasonable granularity.
+Based on our multicriteria selection approach, we selected the configuration ranking second by coherence: n_neighbors = 15, n_components = 10, min_cluster_size = 15, which discovered 7 non-outlier topics.
+This configuration achieved coherence of 0.573 with bootstrap stability of NMI = 0.388 and ARI = 0.164, assigning 23.3% of documents to outliers with the largest topic containing 32.5% of non-outlier documents.
+The selection rationale balanced multiple criteria: while rank 1 achieved marginally higher coherence (0.587), its collapse to only 2 topics was deemed too coarse for meaningful exploration of vaccine perceptions.
+Our selected model provides sufficient granularity (7 topics) to capture distinct thematic dimensions while maintaining strong coherence.
+The moderate stability metrics reflect the broader pattern that high-coherence models show lower NMI due to coarser clustering, consistent with the strong negative correlation documented above.
 4.4 LDA Baseline Comparison
-Systematic grid search across LDA number of topics (5-15 topics) identified 7 topics as optimal. Table 2 compares the optimized LDA baseline with our final BERTopic model.
+We compared our BERTopic model against Latent Dirichlet Allocation using the same coherence-based evaluation framework with a grid search across 3 to 15 topics, and evaluated each configuration using Cv coherence.
+LDA coherence scores ranged from 0.366 (15 topics) to 0.486 (3 topics), with performance generally declining as granularity increased.
+The optimal configuration identified 3 topics with coherence of 0.486.
+Table 3 compares the optimized LDA baseline with our BERTopic model.
 Model Topics Coherence Outliers (%)
-LDA (optimized) 7 0.546 N/A
-BERTopic (merged) 9 0.558 23.6
-Agreement: NMI = 0.122, ARI = 0.080
-Table 2: Comparison of optimized LDA baseline and final hierarchical BERTopic model. Despite comparable coherence, minimal agreement (NMI, ARI) indicates fundamentally different topic structures.
-The minimal agreement between models (NMI = 0.122, ARI = 0.080) indicates that despite comparable coherence values, LDA and BERTopic discover fundamentally different thematic structures.
-This divergence likely reflects architectural differences between LDA’s bag-of-words assumptions and BERTopic’s transformer-based contextual embeddings.
+LDA (optimized) 3 0.486 N/A
+BERTopic (selected) 7 0.573 23.3
+Improvement: +18%
+Table 3: Comparison of optimized LDA baseline and final BERTopic model. BERTopic achieves 18% higher coherence while discovering more than twice as many interpretable topics.
+Our BERTopic model substantially outperformed the baseline, achieving 18% higher coherence (0.573 vs.
+0.486) while discovering more than twice as many interpretable topics (7 vs.
+3).
+This performance gap likely reflects architectural differences: LDA’s bag-of-words assumptions treat documents as unordered token collections, while BERTopic’s transformer-based embeddings capture contextual semantics.
+For conversational data where meaning depends on discourse context (such as “je n’ai pas confiance” referencing different trust domains across utterances) contextualized representations appear better suited to capturing semantic nuance.
+The LDA baseline’s optimal configuration at only 3 topics suggests that without contextualized embeddings, the model struggles to differentiate finer semantic distinctions in French-language focus group discourse.
 4.5 Final Topic Structure and Human Validation
-Human validation by three domain experts confirmed that the 9 merged topics were thematically coherent and interpretable.
-Inter-annotator agreement was robust, with a mean pairwise weighted Cohen’s Kappa of 0.578 (moderate agreement) and a strong mean Spearman’s rank correlation ($\rho$ = 0.725).
-The Intraclass Correlation Coefficient (ICC) for absolute agreement was 0.79 (ICC(2,1)), indicating very good reliability per [koo_guideline_2016].
-Notably, the annotators’ ratings were in perfect agreement within a one-point tolerance (100.0%), underscoring a high level of consensus on topic coherence.
-Table 3 presents the coherence ratings for each topic, along with consensus labels assigned by the annotators.
-The mean coherence scores ranged from 2.00 to 5.00.
-Six of the nine topics received high ratings (mean > 4.0), including those related to Historical trust in national vaccine programs and Polarized attitudes towards official public health sources, both achieving perfect scores of 5.00.
-Conversely, the lowest-rated topic (mean = 2.00) concerned Prevention practices and barriers to care, suggesting that its representative sentences were perceived as less thematically unified.
-There was no significant correlation between a topic’s size (number of sentences) and its coherence score ($\rho$ = -0.213, p = 0.583), indicating that the perceived quality of a topic was independent of its prevalence in the discourse.
+Human validation by three independent domain experts confirmed topic coherence and interpretability.
+The seven emerging themes as labeled by the annotators were: General health knowledge and taboos (Topic 0, 32%), Vaccination importance and institutional processes (Topic 1, 23%), Trusted information sources (Topic 2, 9%), Personal vaccination experiences (Topic 3, 5%), Public versus private healthcare debate (Topic 4, 3%), COVID-19 vaccine adverse effects (Topic 5, 3%), and Institutional trust erosion post-pandemic (Topic 6, 1%).
+Inter-annotator agreement was robust: ICC(2,1) = 0.700 indicates very good reliability per  [koo_guideline_2016], while mean weighted Cohen’s kappa of 0.678 demonstrates substantial agreement [landis1977].
+Annotators achieved perfect agreement within one-point tolerance for 85.7% of topics (6/7).
+Table 4 presents coherence ratings for each topic.
+Five of seven topics received mean ratings $\geq$ 4.0, indicating good-to-excellent coherence.
+Ratings ranged from 1.67 to 4.67 (overall mean = 3.81).
+Topic 0 received the lowest rating (mean = 1.67), reflecting its broad, heterogeneous content spanning multiple health domains.
+Topics 1, 2, 3, 5, and 6 achieved mean ratings between 4.0 and 4.67, confirming strong semantic coherence.
+Topic 4 received moderate ratings (mean = 3.33).
+Topic size showed no correlation with coherence ratings ($\rho$ = -0.213, p = 0.583), confirming evaluations reflected semantic quality rather than prevalence.
 Topic ID Ann. 1 Ann. 2 Ann. 3 Mean Topic Label
-Topic 0 4 4 3 3.67 COVID-19 vaccine side effects and trust erosion
-Topic 1 3 4 3 3.33 Conservative norms and sexual health education
-Topic 2 5 5 5 5.00 Historical trust in national vaccine programs
-Topic 3 4 5 4 4.33 Parental concerns about HPV vaccination
-Topic 4 5 5 4 4.67 Information sources and communication channels
-Topic 5 4 4 5 4.33 Personal experiences with cancer and social impact
-Topic 6 2 2 2 2.00 Prevention practices and barriers to care
-Topic 7 5 5 5 5.00 Polarized attitudes towards official public health sources
-Topic 8 4 5 4 4.33 General disease awareness and knowledge gaps
-Inter-rater reliability: Mean weighted Cohen’s $\kappa$ = 0.578; Mean Spearman’s $\rho$ = 0.725; ICC(2,1) = 0.79
-Table 3: Human validation results for the 9 merged topics. Coherence scores range from 1 (incoherent) to 5 (highly coherent). Representative utterances for each topic are provided in Appendix A.
-The 9 emerging themes range from general awareness and personal experiences with cancer to specific concerns about vaccine side effects and preferred communication channels.
+0 2 2 1 1.67 General health knowledge and taboos
+1 5 5 3 4.33 Vaccination importance and processes
+2 5 5 4 4.67 Trusted information sources
+3 4 5 4 4.33 Personal vaccination experiences
+4 3 4 3 3.33 Public vs. private healthcare debate
+5 4 5 4 4.33 COVID-19 vaccine adverse effects
+6 4 4 4 4.00 Erosion of institutional trust post-COVID
+Inter-rater reliability: Weighted Cohen’s $\kappa$ = 0.678; ICC(2,1) = 0.700
+Table 4: Human validation results. Coherence scores range from 1 (incoherent) to 5 (highly coherent).
 Figure 3 visualizes the semantic organization of the topics in two-dimensional embedding space, revealing distinct thematic clusters and their spatial relationships.
@@
 Each point represents a document (utterance) from the focus group transcripts.
-Colored regions show the 9 merged topics, with spatial proximity indicating semantic similarity.
+Colored regions show the 7 merged topics, with spatial proximity indicating semantic similarity.
 5 Discussion
-5.1 Bootstrap Stability Analysis and the NMI-ARI Divergence
-Our implementation of bootstrap stability analysis builds on Greene et al.
-[greene2014]’s recommendations of performing bootstrap resampling for assessing topic model robustness.
-A critical empirical finding concerns the divergent behavior of NMI and ARI metrics across the parameter space.
-While NMI values ranged from 0.376 to 0.700 with highest stability at smaller cluster sizes (min_cluster_size = 5), ARI showed the opposite pattern with values of 0.094 to 0.205 and peak performance at larger cluster sizes (min_cluster_size = 15).
-This divergence reflects the fundamental differences in what these metrics quantify.
-ARI (Adjusted Rand Index) measures exact agreement in pairwise cluster assignments [hubert1985].
-It is highly sensitive to boundary shifts: when documents near cluster boundaries are reassigned across bootstrap samples, this creates pairwise disagreements that lower ARI scores.
-This sensitivity is particularly pronounced for small clusters, where even minor document reassignments represent large proportional changes to cluster composition.
-In contrast, NMI (Normalized Mutual Information) measures whether clustering solutions preserve the same underlying information structure based on entropy [strehl2002].
-NMI remains relatively stable when fine-grained semantic distinctions are consistently recovered, even if exact cluster boundaries shift somewhat across bootstrap samples.
-Recent systematic evaluations of topic modeling algorithms have demonstrated that ARI and NMI can favor different model configurations depending on data characteristics [rudiger_topic_2022], confirming that neither metric is universally optimal.
-For our exploratory analysis of focus group discussions, where the goal is to discover the full thematic landscape including both dominant and peripheral themes, we argue that NMI provides a more appropriate stability criterion.
-Fine-grained, information-rich models better capture the nuanced structure of focus group discourse, including smaller but important sub-themes that inform health communication interventions.
-5.2 Navigating the Stability-Coherence Tradeoff
-The stability-coherence tradeoff we observed, with smaller cluster sizes yielding higher stability (NMI = 0.695) but lower coherence (0.392), while larger cluster sizes produced higher coherence (0.589) but lower stability (NMI = 0.376), is an example of the multi-dimensional optimization challenge discussed by Hosseiny and Baumer [hosseiny2023], who emphasized that no single optimal configuration exists for all contexts.
-Our hierarchical merging strategy offers a practical solution: by extracting 49 fine-grained topics with high stability then consolidating to 9 super-topics, we achieved coherence (0.558) superior to direct extraction of 9 topics (0.539) while maintaining stability benefits.
-The hierarchical approach also addresses a practical challenge: fine-grained models capture nuanced distinctions but may overwhelm readers, while coarse models sacrifice detail for simplicity.
-By explicitly modeling topic hierarchies, researchers can present high-level summaries while preserving access to fine-grained distinctions for detailed analysis, which is particularly valuable for focus group research where different stakeholders require different levels of granularity.
+5.1 Bootstrap Stability: A Multifaceted Construct
+Our bootstrap stability analysis revealed that partition similarity metrics capture fundamentally different aspects of clustering agreement.
+NMI and ARI exhibited a strong negative correlation (r = -0.691), systematically disagreeing on which configurations represent stable models.
+NMI favored fine-grained configurations (min_cluster_size = 5) producing 34-55 topics with high information structure preservation (NMI = 0.63-0.70), while ARI favored coarser configurations (min_cluster_size = 15) with fewer topics where exact pairwise assignments remained more consistent (ARI = 0.15-0.20).
+This divergence reflects fundamental differences in what these metrics quantify.
+NMI measures whether clustering solutions preserve the same information structure based on entropy, remaining stable when fine-grained semantic distinctions are consistently recovered across bootstrap samples even if exact cluster boundaries shift [strehl2002].
+ARI measures exact pairwise assignment agreement, highly sensitive to boundary variations particularly in small clusters where minor document reassignments represent large proportional changes [hubert1985].
+Recent systematic evaluations have confirmed that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that "stability" is not a monolithic construct but depends on which aspect of partition agreement is prioritized.
+Moreover, these stability metrics showed different relationships with topic interpretability.
+NMI exhibited a strong negative correlation with coherence (r = -0.911), while ARI showed moderate positive correlation (r = 0.639) though with limited discriminative power due to universally low values.
+This metric-dependent relationship demonstrates that bootstrap stability and semantic quality represent distinct dimensions that may conflict rather than align.
+These findings suggest that stability assessment should employ multiple metrics to capture different facets of partition robustness, and that stability measures alone provide insufficient guidance for model selection in exploratory research contexts where interpretability is paramount.
+5.2 Multi-Criteria Model Selection and Validation
+The competing objectives revealed by our systematic evaluation (possible stability-coherence tradeoff, varying topic granularity, and different coverage patterns) demonstrate why rigid optimization strategies prove insufficient in practice.
+While the highest-coherence configuration collapsed to only 2 topics, our selection of 7 topics reflects a judgment that thematic resolution matters alongside semantic quality.
+This decision acknowledges that optimal configurations depend on research goals: classification tasks might prioritize stability for reproducibility, while exploratory analysis requires balancing interpretability with analytical utility.
+Human validation by three domain experts confirmed that coherence-prioritized selection produced interpretable topics, with robust inter-annotator agreement (ICC = 0.700, $\kappa$ = 0.678) and the majority of topics receiving high ratings.
+Importantly, ratings reflected semantic quality rather than topic prevalence, confirming that computational coherence metrics successfully predict expert judgments of interpretability in our application domain.
+The alignment between automated Cv scores and human assessments addresses longstanding concerns that algorithmic optimization may produce statistically optimal but substantively incoherent topics [chang2009].
+Our multi-criteria framework provides a transparent, defensible alternative to single-metric optimization or default configurations.
 5.3 Comparison with LDA Baseline
-We compared our final hierarchical BERTopic model to a systematically tuned LDA baseline.
-Grid search across six configurations identified 7 topics as optimal for LDA, achieving coherence of 0.5464, marginally lower than our BERTopic model (9 topics, coherence = 0.558).
-This small difference suggests comparable performance on semantic coherence.
-However, the minimal agreement between topic assignments (NMI = 0.122, ARI = 0.080) indicates fundamentally different thematic structures.
-This divergence suggests that architectural differences (LDA’s bag-of-words assumptions versus BERTopic’s transformer-based embeddings) lead to qualitatively distinct topic decompositions.
-For conversational data where meaning depends on discourse context (e.g., "je n’ai pas confiance" referencing different trust domains), contextual embeddings may capture nuances that statistical co-occurrence patterns miss.
-The near-independence of topic assignments (NMI = 0.122) suggests these methods may be complementary rather than competitive, capturing different facets of discourse structure.
-Future work with parallel human evaluation of both LDA and BERTopic topics could clarify whether these differences reflect genuinely distinct conceptualizations or alternative representations of similar themes.
-5.4 Human Validation of Topic Quality
-Our human evaluation by three independent domain experts provided essential confirmation that our computationally optimized topics align with substantive interpretations.
-The validation yielded robust inter-annotator agreement across multiple metrics.
-The Intraclass Correlation Coefficient (ICC(2,1) = 0.79) indicates very good reliability [koo_guideline_2016], demonstrating consistent evaluation across raters.
-Mean pairwise weighted Cohen’s kappa of 0.578 (moderate agreement) accounts for the ordinal nature of ratings while showing substantial consensus.
-Most strikingly, annotators achieved perfect agreement within a one-point tolerance (100%), indicating high consistency in perceived topic quality even when exact ratings differed marginally.
-Six of nine topics received high ratings (mean > 4.0), with two achieving perfect scores (5.0).
-Only one topic received a lower rating (mean = 2.0), signaling potential for refinement.
-Topic size showed no correlation with coherence ($\rho$ = -0.213, p = 0.583), confirming that ratings reflected semantic quality rather than prevalence.
-These findings validate that our systematic computational pipeline produces topics that domain experts recognize as interpretable and meaningful, addressing longstanding concerns [nanda2023, steele2025] that topic modeling algorithms may generate statistically optimal but substantively incoherent clusters.
-The strong agreement across three independent raters further suggests that the identified topics capture genuine thematic structures rather than analyst-specific interpretations.
-However, our validation focused exclusively on the final 9 merged topics; the 49 fine-grained initial topics were not independently evaluated.
-Future work could examine whether human raters prefer fine-grained versus consolidated topic structures, and whether stability-optimized configurations yield more interpretable topics than coherence-optimized alternatives.
-5.5 Limitations
+The substantial performance gap between BERTopic and LDA (18% higher coherence, 7 vs.
+3 topics) likely reflects fundamental architectural differences.
+LDA’s bag-of-words assumptions treat documents as unordered token collections, relying on statistical co-occurrence patterns to infer latent topics.
+In contrast, BERTopic’s transformer-based sentence embeddings capture contextual semantics, preserving meaning that depends on word order and discourse context.
+For conversational focus group data, this distinction proves particularly important.
+The same phrase may reference different concepts depending on surrounding discourse.
+Contextual embeddings can distinguish these semantic nuances, while bag-of-words representations collapse them into undifferentiated co-occurrence statistics.
+The LDA baseline’s optimal configuration at only 3 topics suggests that without contextualized representations, the model struggles to differentiate finer semantic distinctions, producing overly broad topics that sacrifice thematic resolution for coherence.
+However, our comparison has limitations.
+Following standard practice [hagg2022], we tuned only the number of topics for LDA while maintaining default hyperparameters for Dirichlet priors and learning parameters.
+More extensive hyperparameter exploration could potentially improve LDA performance, though we consider our approach representative of typical applied usage.
+Additionally, we did not compare against other neural topic modeling approaches such as Top2Vec or CTM, limiting our ability to assess BERTopic’s performance relative to alternative contextualized methods.
+5.4 Limitations and Future Directions
 Several limitations merit consideration.
-First, the substantial sensitivity of results to hyperparameter choices raises questions about generalizability.
-Our finding that min_cluster_size alone shifted stability from 0.376 to 0.700, coherence from 0.385 to 0.589, and topic count from 2 to 52 demonstrates that configuration choices profoundly affect outcomes.
-While our systematic exploration represents a more rigorous approach than typical practice, specific optimal configurations for Tunisian HPV vaccine focus groups may not transfer to focus groups on different topics, with different conversational dynamics, or in different cultural contexts.
-Second, our study is limited to French-language transcripts from Tunisia.
+First, our study is restricted to French-language focus group transcripts from Tunisia.
 The multilingual embedding model we employed may perform differently across languages, and the identified topic structures likely reflect culturally specific patterns of health discourse.
 Validation in other linguistic and cultural contexts is needed to assess the broader applicability of our methodological framework.
-Third, our LDA baseline was tuned only on the number of topics following current practice [hagg2022], potentially underestimating LDA’s optimal performance.
-More extensive tuning of prior distributions (alpha, beta) and learning parameters could improve LDA coherence.
-More critically, we did not conduct parallel human evaluation of LDA topics, which would have provided direct empirical comparison of interpretability beyond algorithmic metrics.
-Finally, we compared BERTopic only with LDA, not with other neural topic modeling approaches such as Top2Vec or CTM (Contextualized Topic Models).
-Egger and Yu [egger2022] found notable performance differences across methods depending on corpus characteristics, suggesting value in comprehensive benchmarking studies.
+Second, the substantial sensitivity of results to hyperparameter choices raises questions about generalizability.
+While our systematic exploration represents a more rigorous approach than accepting default settings, specific optimal configurations may not transfer to focus groups with different conversational dynamics or substantive domains.
+Third, one large topic received poor coherence ratings from evaluators, indicating the model produced a heterogeneous catch-all category rather than a unified theme.
+This suggests limitations in BERTopic’s density-based clustering for highly diverse discourse, where some utterances may not fit coherent thematic patterns.
+Finally, we compared only against LDA rather than other neural methods, and our LDA baseline was tuned only on number of topics following standard practice.
+More comprehensive benchmarking across methods and fuller hyperparameter exploration would strengthen comparative conclusions.
+Future work should examine whether the stability-coherence tradeoff generalizes across languages and data types, explore alternative stability metrics that better align with interpretability, and conduct systematic comparisons with other neural topic modeling approaches.
+Integration with manual qualitative coding frameworks could also clarify how computational topics relate to human-identified themes.
 5.5 Implications for Focus Group Research
-Despite these limitations, our work demonstrates the feasibility and value of applying rigorous topic modeling to focus group transcripts.
-To our knowledge, this is the first study to develop and validate a systematic topic modeling framework specifically for focus group discussions.
-This contribution is meaningful because focus groups remain a cornerstone methodology in qualitative health research, yet the labor-intensive nature of traditional coding limits the scale of synthesis possible.
-Computational approaches do not replace human interpretation but can augment it by efficiently identifying thematic patterns across multiple sessions.
-Our findings suggest practical recommendations for researchers considering topic modeling for focus group analysis.
-First, invest in systematic hyperparameter exploration rather than accepting default configurations.
-Performance varies dramatically, and optimal settings depend on corpus characteristics.
-Second, prioritize stability assessment through bootstrap resampling.
-Coherence scores alone provide incomplete quality information.
-Third, use multiple complementary metrics (stability, coherence, coverage, granularity) and be transparent about tradeoffs in model selection.
-Fourth, validate computational outputs through human review by domain experts.
-Finally, consider hierarchical approaches that extract fine-grained topics for stability then consolidate for interpretability.
-This two-stage strategy can navigate the stability-coherence tradeoff effectively.
+This work demonstrates the feasibility and value of applying systematic topic modeling to focus group transcripts.
+Focus groups remain a cornerstone methodology in qualitative health research, yet the labor-intensive nature of traditional manual coding limits the scale of synthesis possible across multiple sessions.
+Computational approaches do not replace human interpretation but can augment it by efficiently identifying thematic patterns, particularly valuable for large-scale focus group studies where manual analysis of hundreds of transcripts becomes prohibitive.
+Our findings suggest several practical recommendations for researchers considering topic modeling for focus group analysis.
+First, invest in systematic hyperparameter exploration rather than accepting default configurations, as performance varies dramatically depending on corpus characteristics.
+Second, evaluate models across multiple complementary metrics (stability, coherence, granularity, and coverage) and transparently document the tradeoffs inherent in final selection decisions.
+Third, validate computational outputs through independent human review by domain experts using structured protocols; we recommend at least three raters evaluating representative documents per topic on standardized coherence scales.
+Fourth, apply minimal preprocessing for transformer-based models, preserving stop words and contextual elements that bag-of-words approaches would discard.
+These practices balance computational efficiency with the interpretive rigor qualitative research demands, providing a principled framework for augmenting traditional focus group analysis with algorithmic support.
 6 Conclusion
-This paper presents a systematic, reproducible computational framework for applying BERTopic to focus group transcripts, contributing validated methodological practices to the computational social science toolkit.
+This paper presents a systematic, reproducible computational framework for applying BERTopic to focus group transcripts, contributing validated methodological practices to the qualitative research toolkit.
 Through rigorous exploration of 27 hyperparameter configurations with bootstrap stability assessment and formal human validation, we establish that transformer-based topic modeling can reliably extract interpretable themes from structured conversational data.
-The strong inter-rater agreement (ICC = 0.79) confirms that computational approaches, when properly validated, produce outputs aligned with expert domain understanding.
+The strong inter-rater agreement (ICC = 0.700) confirms that computational approaches, when properly validated, produce outputs aligned with expert domain understanding.
 Our work addresses a critical gap: while topic modeling has been extensively applied to social media and online text, focus group discussions, a cornerstone of qualitative research, have received limited computational attention.
-By demonstrating feasibility on transcripts with naturally imbalanced topic distributions and relatively small corpus size (1,076 utterances), we extend neural topic modeling to structured research settings where traditional coding remains dominant.
+By demonstrating feasibility on transcripts with naturally imbalanced topic distributions and relatively small corpus size (1,075 utterances), we extend neural topic modeling to structured research settings where traditional coding remains dominant.
 The transparent documentation of all modeling choices, evaluation procedures, and quality tradeoffs provides a reproducible template that researchers can adapt to their own focus group contexts.
-This methodological foundation enables substantive analysis of HPV vaccine perceptions in Tunisia, where our identified themes reveal specific patterns in health discourse, including vaccine hesitancy drivers, information source credibility, gender-specific concerns, and trust in health institutions.
+This methodological foundation enables substantive analysis of vaccine perceptions in Tunisia, where our identified themes reveal specific patterns in health discourse, including vaccine hesitancy drivers, information source credibility, and trust in health institutions.
 These findings will inform targeted communication strategies and policy interventions in the forthcoming companion paper.
@@
 Appendix A Appendix: Representative Topic Examples
-Table 4 provides representative utterances for each of the 9 merged topics identified by our hierarchical BERTopic model, presented in both original French and English translation.
+Table 5 provides representative utterances for each of the 7 topics identified by our BERTopic model, presented in both original French and English translation.
 These examples illustrate the thematic content and semantic coherence of each topic cluster.
 Topic Label Original (French) Translation (English)
-0 - COVID-19 vaccine side effects and trust erosion Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire après le vaccin contre le Covid.
-Je suis devenu quelqu’un qui oublie beaucoup, et d’autres confirment aussi ce constat.
-Moi, sincèrement, avant la crise du COVID, j’avais une confiance totale dans notre système de santé et je n’ai jamais pensé que c’était nécessaire de poser des questions, surtout que la Tunisie a une longue tradition en matière de vaccination.
-Mais après le COVID, je n’ai plus confiance.
+0 - General disease knowledge, health fears, and taboos Je pense que c’est un sujet délicat et on ne peut pas en parler avec tout le monde.
+Oui, exactement.
+Ma collègue m’a confié que ce qu’elle vivait dépassait ma compréhension.
+Par exemple, les douleurs qu’elle ressentait aux toilettes, les symptômes étranges… C’est une maladie existante mais rare.
+I think this is a sensitive topic, and you cannot talk about it with everyone.
+Yes, exactly.
+My colleague confided that what she was going through was beyond my understanding.
+For example, the pain she felt when using the bathroom, the strange symptoms… It is a real but rare disease.
+1 - General vaccination importance Moi, je suis convaincue qu’il faut faire tous les vaccins, bien que je n’aie pas encore d’enfants.
+La question de la vaccination obligatoire est importante depuis toujours, car les établissements scolaires et d’autres institutions exigent des élèves, étudiants et employés ce document, le certificat de vaccination.
+La majorité d’entre nous suivent le carnet de vaccination.
+I am convinced that one should take all vaccines, even though I do not yet have children.
+The issue of mandatory vaccination has always been important because schools and other institutions require students and employees to provide this document: the vaccination certificate.
+Most of us follow the vaccination schedule.
+2 - Trusted sources of health information Je prends l’information de tout médecin ou spécialiste, qu’il travaille dans le secteur public ou privé.
+Il ne faut pas dépendre des médecins.
+Chacun peut s’informer par ses propres moyens.
+Facebook est une source importante pour moi et j’ai confiance dans les informations publiées sur Facebook.
+Je ne suis pas d’accord avec elle.
+Le ministère de la Santé n’est pas n’importe qui.
+Il faut suivre les instructions du ministère.
+I take information from any doctor or specialist, whether they work in the public or the private sector.
+You should not depend on doctors.
+Everyone can inform themselves on their own.
+Facebook is an important source for me, and I trust the information published there.
+I do not agree with her.
+The Ministry of Health is not just anyone.
+You must follow the ministry’s instructions.
+3 - Personal experiences with vaccination Moi, je ne l’ai pas reçu car j’étais enceinte à l’époque de la vaccination et j’ai attrapé le virus.
+J’ai beaucoup souffert et j’ai eu de gros problèmes respiratoires.
+Je n’étais pas convaincue, mais j’ai dû me faire vacciner.
 Le vaccin contre la grippe, je l’ai fait, c’est un peu cher, cette année je ne l’ai pas fait.
-Sorry, but based on my personal experience and what I hear, the majority of people have had memory problems after the Covid vaccine.
-I have become someone who forgets a lot, and others also confirm this observation.
-Honestly, before the COVID crisis, I had complete trust in our healthcare system and I never thought it was necessary to ask questions, especially since Tunisia has a long tradition of vaccination.
-But after COVID, I no longer have trust.
-I got the flu vaccine; it’s a bit expensive.
-This year, I didn’t get it.
-1 - Conservative norms and sexual health education Vous savez, la société tunisienne est conservatrice, même lorsque quelque chose comme ça se produit, on n’en parle pas, c’est honteux.
-Parce que les filles sont devenues plus libres, nous avons fait une campagne de sensibilisation dans les lycées pour expliquer aux filles et aux garçons l’importance de prendre un préservatif pour se protéger.
-Mais, nous avons été stigmatisés.
-Les gens ont dit que nous sommes en train d’encourager les filles à être encore plus libres.
-Alors que c’est juste de la prévention.
-Non, en 4ème, par exemple, on lui dit que le couple est marié et que la mère est enceinte.
-Mais cela n’explique pas l’acte sexuel.
-You know, Tunisian society is conservative; even when something like that happens, people don’t talk about it, it’s shameful.
-Because girls have become freer, we conducted an awareness campaign in high schools to explain to girls and boys the importance of taking a condom to protect themselves.
-But we were stigmatized.
-People said that we were encouraging girls to be even freer.
-Whereas it is just prevention.
-No, in 8th grade, for example, they are told that the couple is married and that the mother is pregnant.
-But that does not explain the sexual act.
-2 - Historical trust in national vaccine programs Les vaccins sont considérés comme un programme réussi.
-Lorsque vous comparez la Tunisie à des pays d’Afrique où ils sont confrontés à de nombreuses maladies graves en raison du manque de vaccins nécessaires, nous devons apprécier ce que nous avons.
-La vaccination en général est une bonne chose car elle protège contre la poliomyélite, l’hépatite et ceux qui ne reçoivent pas les vaccinations nécessaires peuvent avoir des problèmes de santé plus tard.
-Dans le public, ils ont leur carnet de vaccination.
-Avant de rentrer à l’école, ils doivent être vaccinés.
-Vaccines are considered a successful program.
-When you compare Tunisia to African countries where they face many serious diseases due to the lack of necessary vaccines, we must appreciate what we have.
-Vaccination in general is a good thing because it protects against poliomyelitis, hepatitis, and those who do not receive the necessary vaccinations may have health problems later on.
-In public, they have their vaccination record.
-Before entering school, they must be vaccinated.
-3 - Parental concerns about HPV vaccination Principalement pour protéger la santé de ma fille.
-Comme c’est la cadette, je voulais la protéger autant que possible.
-Quand tu lui dis que la maladie est liée au col de l’utérus et qu’elle ne concerne que les filles, certainement que ces données le feront hésiter et penser que le vaccin pourrait peut-être affecter la fertilité de sa fille.
-C’est pourquoi j’insiste sur l’importance de mener des campagnes de sensibilisation pour expliquer aux gens que les filles représentent la population la plus touchée par ce virus.
-Mainly to protect my daughter’s health.
-Since she is the youngest, I wanted to protect her as much as possible.
-When you tell him that the disease is related to the cervix and only affects girls, these facts will certainly make him hesitate and think that the vaccine might possibly affect his daughter’s fertility.
-That is why I emphasize the importance of conducting awareness campaigns to explain to people that girls are the population most affected by this virus.
-4 - Information sources and communication channels Les médecins doivent être formés à communiquer efficacement avec les parents.
-La télévision atteint un plus grand nombre de citoyens et on ne peut pas mentir à la télévision.
-Les gens dans les réseaux sociaux numériques publient des informations fausses et les gens finissent par les croire.
-Par exemple, récemment, ils ont dit qu’il y avait de la neige à Ain Draham alors qu’il n’y avait ni neige ni rien.
-C’est de la désinformation.
-Doctors must be trained to communicate effectively with parents.
-Television reaches a larger number of citizens and one cannot lie on television.
-People on digital social networks publish false information and people end up believing it.
-For example, recently, they said there was snow in Ain Draham when there was neither snow nor anything.
-This is misinformation.
-5 - Personal experiences with cancer and its social impact Je connais quelqu’un qui en a souffert et qui a tout enlevé, elle avait 38 ans, elle en a été affectée et déprimée.
-Elle est mariée et a un enfant, c’est ma cousine, elle a été hospitalisée, elle a subi une opération pour enlever son utérus.
-Il y a aussi une maladie très répandue qui est le cancer du foie.
-J’ai subi deux opérations chirurgicales pour l’enlever.
-N’y a-t-il pas de vaccination contre cela ?
-J’ai une voisine qui est tombée malade de cette maladie et son mari l’a divorcée.
-Les hommes ne soutiennent pas les femmes dans ces situations.
-I know someone who suffered from it and had everything removed; she was 38 years old, and it affected her deeply and made her depressed.
-She is married and has a child; she is my cousin.
-She was hospitalized and underwent surgery to have her uterus removed.
-There is also a very common disease which is liver cancer.
-I have undergone two surgical operations to remove it.
-Is there no vaccination against it?
-I have a neighbor who fell ill with this disease and her husband divorced her.
-Men do not support women in these situations.
-6 - Prevention practices and barriers to care Moi, je ne veux pas faire le frottis car j’ai peur.
-J’ai eu une mauvaise expérience après mon accouchement dans un hôpital public et je ne veux pas que quiconque me touche dans ma zone intime.
-Oui, je consulte mon médecin chaque année pour un contrôle.
-Il y a des remèdes naturels que j’utilise quand mon enfant est malade, comme le miel et l’eucalyptus.
-I don’t want to have the smear test because I’m afraid.
-I had a bad experience after my delivery in a public hospital and I don’t want anyone to touch me in my intimate area.
-Yes, I see my doctor every year for a check-up.
-There are natural remedies that I use when my child is sick, such as honey and eucalyptus.
-7 - Polarized attitudes towards official public health sources Je ne sais pas, je n’ai pas confiance dans l’ensemble du système.
-Nous faisons confiance aux vaccins approuvés par l’État et le ministère de la Santé.
-Moi, je fais confiance à la page officielle du ministère de la Santé.
-I don’t know, I don’t trust the whole system.
-We trust vaccines approved by the state and the Ministry of Health.
-I trust the official page of the Ministry of Health.
-8 - General disease awareness and knowledge gaps C’est la première fois que j’entends parler de cette maladie.
-L’hygiène est l’une des principales causes de cette maladie.
-Je n’ai entendu parler d’aucun cas, je sais seulement que c’est une maladie maligne.
-This is the first time I have heard of this disease.
-Hygiene is one of the main causes of this disease.
-I have not heard of any cases; I only know that it is a malignant disease.
-Table 4: Representative utterances for each of the 9 merged topics with original French text and English translations.
+I did not receive it because I was pregnant at the time of vaccination, and I caught the virus.
+I suffered a lot and had severe respiratory problems.
+I was not convinced, but I had to get vaccinated.
+I got the flu vaccine; it is a bit expensive.
+This year, I did not get it.
+4 - Private vs Public sector debate C’est injuste.
+Vous pensez que c’est normal que je paie cette somme pour un vaccin que j’ai normalement le droit d’obtenir gratuitement !
+C’est mon droit !
+Mais il y a plusieurs lacunes dans le secteur public qu’il faut reconnaître.
+Le secteur privé dans le domaine de la santé est devenu supérieur au secteur public… Parce que ma fille a suivi sa grossesse chez un médecin privé et a accouché dans une clinique privée.
+Elle n’a donc pas eu recours au secteur public dès le début, choisissant de surveiller son enfant chez un pédiatre privé.
+It is unfair.
+Do you think it is normal for me to pay this amount for a vaccine that I have the right to obtain for free?
+It is my right!
+But several shortcomings in the public sector must be acknowledged.
+The private health sector has become superior to the public sector… Because my daughter followed her pregnancy with a private doctor and gave birth in a private clinic.
+She therefore did not use the public sector from the beginning, choosing to monitor her child with a private pediatrician.
+5 - COVID-19 vaccine adverse effects Le Covid a vraiment changé la donne.
+Personnellement, après avoir reçu le vaccin contre le Covid, mon immunité a diminué.
+Avant, je guérissais d’une grippe en une semaine, mais maintenant, cela prend au moins un mois.
+Moi, mon ophtalmologue m’a critiqué car j’ai fait le vaccin contre le Covid.
+J’ai eu des problèmes dans mes yeux après avoir fait ce vaccin.
+Le vaccin contre le COVID provoque Alzheimer.
+Covid really changed things.
+Personally, after receiving the Covid vaccine, my immunity decreased.
+Before, I recovered from the flu in one week, but now it takes at least a month.
+My ophthalmologist criticized me for taking the Covid vaccine.
+I had eye problems after receiving it.
+The Covid vaccine causes Alzheimer’s.
+6 - Erosion of institutional trust post pandemic Après la pandémie de COVID, l’opinion publique est devenue plus prudente.
+Non, auparavant, nous avions une confiance aveugle dans le ministère de la Santé, mais les failles sont apparues après la pandémie de COVID, et notre confiance dans le ministère de la Santé a diminué… Le COVID a créé ce problème, avant il n’y avait pas de précautions contre ceci ou cela.
+Depuis que nous avons appris les risques du vaccin et qu’il peut provoquer une crise cardiaque, les gens ont peur.
+After the COVID pandemic, public opinion became more cautious.
+No, previously we had blind trust in the Ministry of Health, but flaws became apparent after the COVID pandemic, and our trust in the Ministry of Health decreased… COVID created this problem; before, there were no precautions about this or that.
+Since we learned about the risks of the vaccine and that it can cause a heart attack, people are afraid.
+Table 5: Representative utterances for topics with original French text and English translations.